### v5

# Enumeration Kernels of Polynomial Size for Cuts of Bounded Degree

Abstract
Enumeration kernelization was first proposed by Creignou et al. [TOCS 2017] and was later refined by Golovach et al. [JCSS 2022] into two different variants: fully-polynomial enumeration kernelization and polynomial-delay enumeration kernelization. In this paper, we consider the Degree-$d$-Cut problem from the perspective of (polynomial-delay) enumeration kenrelization. Given an undirected graph $G=(V,E)$, a cut $F=E(A,B)$ is a degree-$d$-cut of $G$ if every $u\in A$ has at most $d$ neighbors in $B$ and every $v\in B$ has at most $d$ neighbors in $A$. Checking the existence of a degree-$d$-cut in a graph is a well-known NP-Complete problem and is well-studied in parameterized complexity [Algorithmica 2021, IWOCA 2021]. This problem also generalizes a well-studied problem Matching Cut (set $d=1$) that has been a central problem in the literature of polynomial-delay enumeration kernelization. In this paper, we study three different enumeration variants of this problem, Enum Deg-$d$-Cut, Enum Min Deg-$d$-Cut and Enum Max Deg-$d$-Cut that intends to enumerate all the degree-$d$-cuts, all the minimal degree-$d$-cuts and all the maximal degree-$d$-cuts respectively. We consider various structural parameters of the input and provide polynomial-delay enumeration kernels for Enum Deg-$d$-Cut and Enum Max Deg-$d$-Cut and fully-polynomial enumeration kernels of polynomial size for Enum Min Deg-$d$-Cut.
1 Introduction
Given a simple graph $G=(V,E)$, a cut is a bipartition of its vertex set $V(G)$ into two nonempty sets $A$ and $B$, denoted by $(A,B)$. The set of all edges with one endpoint in $A$ and the other in $B$ is called the edge cut or the set of crossing edges of $(A,B)$. We use $F=E_{G}(A,B)$ to denote the edge cut of $(A,B)$. If the graph is clear from the context, we omit the subscript. A cut $(A,B)$ of $G$ is said to be a matching cut if every vertex of $A$ has at most one neighbor in $B$ and every vertex of $B$ has at most one neighbor in $A$. Alternatively, a cut $(A,B)$ is a matching cut if the set of crossing edges of $(A,B)$ is a matching. The Matching Cut problem asks whether a given graph has a matching cut. Gomes and Sau [15] introduced a natural generalization of matching cut, called $d$-cut which is defined as follows: a cut $(A,B)$ is a $d$-cut if every vertex of $A$ has at most $d$ neighbors in $B$ and every vertex of $B$ has at most $d$ neighbors in $A$. Note that a $1$-cut is a matching cut. Moreover, not every graph admits a $d$-cut. For a fixed integer $d\geq 1$, the Degree-$d$-Cut problem (decision version) asks if an undirected graph $G$ has a $d$-cut. If we put $d=1$, then the Degree-$d$-Cut problem is the same as the Matching Cut. Both Matching Cut and Degree-$d$-Cut are known to be NP-Complete problems even in special classes of graphs [8, 9, 6, 22, 24, 25, 28, 30] and have been explored from the perspective of parameterized complexity [21, 15, 1, 3, 2].
Kernelization [13, 17, 19] is a central subfield of parameterized complexity that was initially explored for decision problems. Informally, the objective of a kernelization algorithm (or kernel) is to shrink the input instance into an equivalent instance by a polynomial-time procedure the size of which is dependent only on the parameter. The efficiency of a kernelization (or a preprocessing algorithm) is determined by the size bound of the kernel. Later, several extensions of kernelization have been adopted for parameterized optimization problems [27], counting problems [26], and enumeration problems [14, 12, 20]. In the case of parameterized enumeration problems, a small kernel can be viewed as a “compact representation” of all the feasible solutions of the original instance. The first attempt was suggested by Damaschke [14] where the aim was to shrink the input instance to an output instance containing all minimal solutions of size at most $k$. But, this definition was limited to only minimization problems. Being promoted by this drawback, Creignou et al. [11, 12, 10] found some weaknesses and introduced a modified notion in which they had asked for two algorithms. The first algorithm is kernelization algorithm that shrinks the input instance to an output instance (the kernel) the size of which is bounded by the parameter. The second algorithm is solution-lifting algorithm that given a solution of the kernel, outputs a collection of solutions to the input instance. But this notion allows the solution-lifting algorithm to run in FPT-delay. In such a case, the solution-lifting algorithm can output all solutions to the input instance by ignoring the kernel. This prompted Golovach et al. [20] to introduce two refined notions of enumeration kernel: (i) fully-polynomial enumeration kernel and (ii) polynomial-delay enumeration kernel: in which the solution-lifting algorithms must run in polynomial-time and polynomial-delay respectively (see Definitions 2.1). It has been justified [20] that this definition has a desirable consequence: (i) an enumeration problem admits a polynomial-delay (or polynomial-time) enumeration algorithm if and only if it admits a polynomial-delay (fully-polynomial) enumeration kernel of constant size. This is analogous to the fact that a problem admits a polynomial-time algorithm if and only if it admits a kernel of constant size.
Additionally, this above definition [20] also ensures us that a parameterized enumeration problem admits an FPT-delay algorithm if and only if it admits a polynomial-delay enumeration kernel. But this automatically enumeration kernel has exponential size with respect to the parameter. This motivates us the following question from the theoretical perspective: if a parameterized enumeration problem admits an FPT-delay algorithm, does it admit a polynomial-delay enumeration kernel of polynomial size? In short, a polynomial-delay enumeration kernel of polynomial size give us a compact representation of all solutions of the original instance, and a guarantee that a collection of ‘equivalent’ solutions can be enumerated with polynomial-delay.
Being motivated by the works on polynomial-delay enumeration kernelization on Matching Cut by Golovach et al. [20] and the works on Degree-$d$-Cut (for every fixed integer $d\geq 1$) by Gomes and Sau [15], we initiate a systematic study on the Degree-$d$-Cut problem from the perspective of enumeration kernelizations; i.e. fully-polynomial enumeration kernelization and polynomial-delay enumeration kernelization. A $d$-cut $F$ is minimal if no proper subset of $F$ is a $d$-cut of $G$ (i.e. inclusion-wise minimal). Similarly, a $d$-cut of $G$ is maximal if no proper superset of $F$ is a $d$-cut of $G$. In this paper, we focus on the following three enumeration variants of Degree-$d$-Cut the formal definition of which are given below.
Enum Deg-$d$-Cut Input: A simple undirected graph $G=(V,E)$. Goal: Enumerate all the $d$-cuts of $G$.
Enum Min Deg-$d$-Cut Input: A simple undirected graph $G=(V,E)$. Goal: Enumerate all minimal $d$-cuts of $G$.
Enum Max Deg-$d$-Cut Input: A simple undirected graph $G=(V,E)$. Goal: Enumerate all maximal $d$-cuts of $G$.
Our Contributions and Motivations Behind the Choice of Parameters:
In this paper, we apply both types of enumeration kernelizations to Enum Min Deg-$d$-Cut, Enum Deg-$d$-Cut, and Enum Max Deg-$d$-Cut. As the decision version Degree-$d$-Cut is NP-hard [9], it is unlikely that Enum Deg-$d$-Cut and other variants admit polynomial-delay enumeration algorithms. The structural parameters that we choose in this paper are the size of a minimum vertex cover (vc), the neighborhood diversity (nd), and the clique-partition number (pc) of the input graph. The motivation behind the choice of these parameters is the following. In the context of a parameterized problem, if it is established that the problem does not allow for a (polynomial-delay) enumeration kernelization of polynomial size concerning a specific parameter, it is logical to opt for a parameter that is either structurally larger or incomparable to the aforementioned parameter. Similarly, if a parameterized problem is acknowledged to permit a (polynomial-delay) enumeration kernelization of polynomial size concerning a particular parameter, it is intuitive/natural to select a parameter that is either structurally smaller or incomparable to the mentioned parameter. None of these problems admit enumeration kernelizations of polynomial size when parameterized by a combined parameter of solution size, maximum degree, and treewidth of the input graph (see [15]). Therefore, it is natural to select minimum vertex cover size (vc) and neighborhood diversity (nd) as parameters that are structurally larger than treewidth and cliquewidth respectively.
Figure 1: An overview of parameter hierarchy. Some relation between the parameters are ${\textsf{pc}}(G)\leq{\sf dtc}(G)+1$, ${\sf cw}(G)\leq{\sf mw}(G)+2$ and ${\sf cw}(G)\leq{\sf cvd}(G)+3$. For every other arrow from parameter $X$ to parameter $Y$ indicates that $X\leq Y$. A green box says the existence of (polynomial-delay) enumeration kernelizations of polynomial size. An orange box indicates non-existence of enumeration kernelizations of polynomial sized unless ${\sf NP\subseteq coNP/poly}$. A pink box indicates the NP-hardness for constant value of parameter. A yellow box says open status in the existence of (polynomial-delay) enumeration kernelizations of polynomial size. The parameters that are marked as $\star$ have been considered in this paper.
We refer to Figure 1 for an illustration of parameters. Additionally, we refer to Table 1 for a short summary of our results.
Theorem 1.1.
For every fixed $d\geq 1$, we prove the following two results.
  1. 1.
    Enum Min Deg-$d$-Cut parameterized by vc admits a fully-polynomial enumeration kernel containing $\mathcal{O}(d^{3}{\textsf{vc}}^{d+1})$ vertices.
  2. 2.
    Enum Deg-$d$-Cut and Enum Max Deg-$d$-Cut parameterized by vc admit polynomial-delay enumeration kernel with $\mathcal{O}(d^{3}{\textsf{vc}}^{d+1})$ vertices.
Next we consider the neighborhood diversity, i.e. nd of the input graph as the parameter and we prove the following result in Section 4.
Theorem 1.2.
For every fixed $d\geq 1$, Enum Min Deg-$d$-Cut parameterized by nd admit a fully-polynomial enumeration kernelization with $\mathcal{O}(d^{2}{\textsf{nd}})$ vertices. Moreover, Enum Deg-$d$-Cut and Enum Max Deg-$d$-Cut admit polynomial-delay enumeration kernelizations with $\mathcal{O}(d^{2}{\textsf{nd}})$ vertices.
Finally, in Section 5, we consider the of a clique partition number, i.e. pc of the input graph as the parameter and prove the following result.
Theorem 1.3.
For every fixed $d\geq 1$, Enum Deg-$d$-Cut, Enum Min Deg-$d$-Cut, and Enum Max Deg-$d$-Cut parameterized by pc admit bijective enumeration kernelizations with $\mathcal{O}({\textsf{pc}}^{d+2})$ vertices.
Parameter $k$ Enum Deg-$d$-Cut Enum Min Deg-$d$-Cut Enum Max Deg-$d$-Cut
[…]
distance to clique $\mathcal{O}({k}^{{d}+2})$-vertex bijective enum-kernel $\mathcal{O}({k}^{{d}+2})$-vertex bi-enum-kernel $\mathcal{O}({k}^{d+2})$-vertex bijective enum-kernel
treewidth, and cliquewidth No polynomial size delay enum-kernel ($\star$) (Proposition 2.2) No polynomial size delay enum-kernel ($\star$) (Proposition 2.2) No polynomial size delay enum-kernel ($\star$) (Proposition 2.2)
[…]
Related Work:
Xiao and Nagamochi [34] had studied a different variant of Degree-$d$-Cut from the perspective of parameterized complexity. We refer to a survey by Wasa [33] for a detailed overview of enumeration algorithms. Bentert et al. [4] have introduced a notion of advice enumeration kernels where they provide an extension to the enumeration problem by Creignou et al. [12] where a small amount of advice is needed rather than the whole input. Counting problems are closely related to enumeration problems. A concrete survey by Thilikos [32] is also available for compact representation of parameterized counting problems. Later, several other studies have also happened in parameterized counting problems, including a few in temporal graph models also (see [29, 18, 7]). Very recently, Lokshtanov et al.[26] have formally introduced the framework of kernelization for counting problems.
2 Preliminaries
Sets, Numbers, and Graph Theory: We use $[r]$ to denote $\{1,\ldots,r\}$ for some $r\in{\mathbb{N}}$ and $A\uplus B$ to denote the disjoint union of the sets $A$ and $B$. For a set $A$, we use $2^{A}$ to denote the power set of $A$. For a finite set $A$ and a nonnegative integer $d\geq 0$, we use (${{A}\choose{d}},{{A}\choose{\leq d}}$, and (${{A}\choose{\geq d}}$ to denote the collection of all subsets of $A$ of size equal to $d$, at most $d$, and at least $d$ respectively. We use standard graph theoretic notations from the book of Diestel [16]. Throughout this paper, we consider simple undirected graphs. For a subset $X\subseteq V(G)$, $G[X]$ denotes the subgraph induced by the vertex subset $X$. Similarly, $G-X$ denotes the graph obtained after deleting the vertex set $X$; i.e. $G-X=G[V(G)\setminus X]$; for a vertex $u$, we use $G-u$ to represent $G-\{u\}$ for the simplicity. Similarly, for a set of edges $F\subseteq E(G)$ (respectively for $e\in E(G)$), we write $G-F$ (respectively $G-e$) to denote the graph obtained after deleting the edges of $F$ (respectively after deleting $e$) from $G$. Equivalently, for an edge $F\subseteq E(G)$, we use $G-F$ to denote the subgraph $G(V,E\setminus F)$ For a vertex $v$, we denote by $N_{G}(v)$ the open neighborhood of $v$, i.e. the set of vertices that are adjacent to $v$ and by $N_{G}[v]$ the closed neighborhood of $v$. Formally, $N_{G}[v]=N_{G}(v)\cup\{v\}$. We omit these subscripts when the graph $G$ is clear from the context. A set of vertices $S\subseteq V(G)$ is called vertex cover if for every $uv\in E(G)$, $u\in S$ or $v\in S$ (or both). The size of a minimum vertex cover of a graph is known as the vertex cover number and is denoted by vc. Given a graph $G=(V,E)$, a set $M\subseteq V(G)$ is said to be a module of $G$ if every pair of vertices of $M$ has the same neighborhood outside $M$. Formally, for every $u,v\in M$, $N_{G}(u)\setminus M=N_{G}(v)\setminus M$. A modular decomposition of $G$ is a decomposition of the vertices of $G$ into vertex-disjoint modules. Formally, ${\mathcal{U}}=\{X_{1},\ldots,X_{k}\}$ is a modular decomposition of $G$ if $\cup_{i=1}^{k}X_{i}=V(G)$ and $X_{i}\cap X_{j}=\emptyset$ for every $i\neq j$. Given a graph $G$, a neighborhood decomposition of $G$ is a modular decomposition ${\mathcal{U}}=\{X_{1},\ldots,X_{k}\}$ of $G$ such that every module is either clique or an independent set. If ${\mathcal{U}}=\{X_{1},\ldots,X_{k}\}$ is a neighborhood decomposition of $G$ with minimum number of modules, then the neighborhood diversity of $G$ is $k$. Such a neighborhood decomposition is also called a minimum-size neighborhood decomposition of $G$. A partition ${\mathcal{C}}=\{C_{1},\ldots,C_{k}\}$ of a graph $G$ is said to be a clique partition of $G$ if for all $1\leq i\leq k$, $C_{i}$ is a clique. The clique partition number of $G$ is denoted by ${\textsf{pc}}(G)$ is the minimum $k$ such that $G$ has a clique partition with $k$ cliques. We introduce a graph operation that we were unable to find in the literature. Let $G$ be an undirected graph such that $u\in V(G)$. We define a graph operation attach a pendant clique $C$ of $r$ vertices into $u$ when a clique $C$ containing $r$ vertices is attached to $G$ such that $C\cap V(G)=\{u\}$. In short form, we often call this graph operation attach an $r$-vertex pendant clique into $u$. The following lemma gives fundamental structural characterizations of a $d$-cuts of $G$.
Lemma 2.1.
Let $G$ be an undirected graph and $(A,B)$ be a $d$-cut of $G$.
  1. 1.
    If $C$ is a clique with at least $12d+$ 1 vertices, then either $C\subseteq A$ or $C\subseteq B$.
  2. 2.
    Let $T$ be a nonempty, proper subset of $V(G)$ such that there are more than $2d$ vertices in $V(G)\setminus T$ whose neighborhoods contain $T$. Then, either $T\subseteq A$ or $T\subseteq B$.
Proof.
We prove the items in the given order.
  1. 1.
    Since $C$ is a clique in $G$ with at least $12d+$ 1 vertices, if $C\cap A,C\cap B\neq\emptyset$, then by pigeon hole principle, either $|C\cap A|\geq d+1$ or $|C\cap B|\geq d+1$. As $C$ is a clique, if $|C\cap A|\geq d+1$, then there is $u\in C\cap B$ that has at least $d+1$ neighbors in $A$. Simiarly, if $|C\cap B|\geq d+1$, then there is $v\in C\cap A$ that has at least $d+1$ neighbors in $A$. Both cases lead to a contradiction that $(A,B)$ is a $d$-cut.
  2. 2.
    If $|T|=1$, then the statement is trivially true. So, we give the proof for nonempty sets $T$ with at least two vertices. For the sake of contradiction, suppose that the premise holds true, but there is a $d$-cut $(A,B)$ of $G$ such that $T\cap A,T\cap B\neq\emptyset$. We consider $v\in V(G)\setminus T$ such that $T\subseteq N_{G}(v)$ and let $I=V(G)\setminus T$. As per the premise, there are at least $12d+$ 1 vertices in $I$ whose neighborhood contains $T$. Let $\widehat{I}_{T}$ be the set of any $12d+$ 1 vertices from $I$ whose neighborhoods contain $T$. That is, for every $v\in\widehat{I}_{T}$, $T\subseteq N_{G}(v)$ and $|\widehat{I}_{T}|\geq 2d+1$. By pigeon hole principle, at least $d+1$ vertices from $\widehat{I}_{T}$ are in $A$ (or in $B$ respectively). If $d+1$ vertices from $\widehat{I}_{T}$ are in $A$, then as $T\cap B\neq\emptyset$, there is a vertex $x\in T\cap B$ such that $x$ has $d+1$ neighbors in $A$. Similarly, if $d+1$ vertices from $\widehat{I}_{T}$ are in $B$, then as $T\cap A\neq\emptyset$, there is a vertex $x\in T\cap A$ such that $x$ has $d+1$ neighbors in $B$. Both the situations lead to a contradiction to the assumption that $(A,B)$ is a $d$-cut of $G$. Similarly, if $d+1$ vertices from $\widehat{I}_{T}$ are in $B$, then there is a vertex $x\in T\cap A$ that has $d+1$ neighbors in $B$. This also contradicts the assumption that $(A,B)$ is a $d$-cut of $G$. Hence, the claim is true.
This completes the proof. ∎
If we put $|T|\geq d+1$ in statement (ii) above, then we get Observation 1 from [15]. We refer to the appendix for the definitions of parameterized complexity due to lack of space.
Parameterized Complexity and Kernelization:
A parameterized problem is denoted by $L\subseteq\Sigma^{*}\times{\mathbb{N}}$ where $\Sigma$ is a finite alphabet. An instance of a parameterized problem is denoted $(x,k)\in\Sigma^{*}\times{\mathbb{N}}$ where and $k$ is considered as the parameter. A parameterized problem $L$ is said to be fixed-parameter tractable if there is an algorithm ${\mathcal{A}}$ that takes an input instance $(x,k)\in\Sigma^{*}\times{\mathbb{N}}$, runs in $f(k)|x|^{\mathcal{O}(1)}$-time and correctly decides if $(x,k)\in L$ for a computable function $f:{\mathbb{N}}\rightarrow{\mathbb{N}}$. The algorithm ${\mathcal{A}}$ is called fixed-parameter algorithm (or FPT algorithm). An important step to design parameterized algorithms is kernelization (or parameterized preprocessing). Formally, a parameterizezd problem $L\subseteq\Sigma^{*}\times{\mathbb{N}}$ is said to admit a kernelization if there is a polynomial-time procedure that takes $(x,k)\in\Sigma^{*}\times{\mathbb{N}}$ and outputs an instance $(x^{\prime},k^{\prime})$ such that (i) $(x,k)\in L$ if and only if $(x^{\prime},k^{\prime})\in L$, and (ii) $|x^{\prime}|+k^{\prime}\leq g(k)$ for some computable function $g:{\mathbb{N}}\rightarrow{\mathbb{N}}$. It is well-known that a decidable parameterized problem is FPT if and only if it admits a kernelization [13].
Parameterized Enumeration and Enumeration Kernelization:
We use the framework for parameterized enumeration was proposed by Creignou et al. [12]. An enumeration problem over a finite alphabet $\Sigma$ is a tuple $(L,{\sf Sol})$ such that
  1. (i)
    $L$ is a decidable problem, and
  2. (ii)
    ${\sf Sol}:\Sigma^{*}\rightarrow 2^{\Sigma^{*}}$ is a computable function such that ${\sf Sol}(x)$ is a nonempty finite set if and only if $x\in L$.
Here $x\in\Sigma^{*}$ is an instance, and ${\sf Sol}(x)$ is the set of solutions to instance $x$. A parameterized enumeration problem is defined as a triple $\Pi=(L,\sf Sol,\kappa)$ such that $(L,{\sf Sol})$ satisfy the same as defined in item (i) and (ii) above. In addition to that $\kappa:\Sigma^{*}\rightarrow{\mathbb{N}}$ is the parameter. We define here the parameter as a function $\kappa(x)$ to the input instance, but it is natural to assume that the parameter is given with the input or $\kappa(x)$ can be computed in polynomial-time. An enumeration algorithm ${\mathcal{A}}$ for a parameterized enumeration problem $\Pi$ is a deterministic algorithm that given $x\in\Sigma^{*}$, outputs ${\sf Sol}(x)$ exactly without duplicates and terminates after a finite number of steps. If ${\mathcal{A}}$ outputs exactly ${\sf Sol}(x)$ without duplicates and eventually terminates in $f(\kappa(x))|x|^{\mathcal{O}(1)}$-time, then ${\mathcal{A}}$ is called an FPT-enumeration algorithm. For $x\in L$ and $1\leq i\leq|{\sf Sol}(x)|$, the $i$-th delay of ${\mathcal{A}}$ is the time taken between outputting the $i$-th and $(i+1)$-th solution of ${\sf Sol}(x)$. The 00-th delay is the precalculation time that is the time from the start of the algorithm until the first output. The $|{\sf Sol}(x)|$-th delay is the postcalculation time that is the time from the last output to the termination of ${\mathcal{A}}$. If the enumeration algorithm ${\mathcal{A}}$ on input $x\in\Sigma^{*}$, outputs ${\sf Sol}(x)$ exactly without duplicates such that every delay is $f(\kappa(x))|x|^{\mathcal{O}(1)}$, then ${{\mathcal{A}}}$ is called an FPT-delay enumeration algorithm.
Definition 2.1.
Let $\Pi=(L,{\sf Sol},\kappa)$ be a parameterized enumeration problem. A polynomial-delay enumeration kernel(ization) for $\Pi$ is a pair of algorithms ${\mathcal{A}}$ and ${\mathcal{A}}^{\prime}$ with the following properties.
  1. (i)
    For every instance $x$ of $\Pi$, the kernelization algorithm ${\mathcal{A}}$ computes in time polynomial in $|x|+\kappa(x)$ an instance $y$ of $\Pi$ such that $|y|+\kappa(y)\leq f(\kappa(x))$ for a computable function $f(\cdot)$.
  2. (ii)
    For every $s\in{\sf Sol}(y)$, the solution-lifting algorithm ${\mathcal{A}}^{\prime}$ computes with delay in polynomial in $|x|+|y|+\kappa(x)+\kappa(y)$ a nonempty set of solutions $S_{s}\subseteq{\sf Sol}(x)$ such that $\{S_{s}\mid s\in{\sf Sol}(y)\}$ is a partition of ${\sf Sol}(x)$.
Observe that by (ii) of the definition above, $x\in L$ if and only if $y\in L$. In the definition above, $|y|+\kappa(y)$ is upper bounded by $f(\kappa(x))$ for some computable function $f:{\mathbb{N}}\rightarrow{\mathbb{N}}$. This computable function $f(\cdot)$ is the size of the kernel. If $f(\cdot)$ is a polynomial function, then $\Pi$ is said to admit a polynomial-delay enumeration kernelization of polynomial size. In addition, the Property (ii) clearly implies that ${\sf Sol}(x)\neq\emptyset$ if and only if ${\sf Sol}(y)\neq\emptyset$. An enumeration kernel is bijective if for every $s\in{\sf Sol}(y)$, the solution-lifting algorithm produces a unique solution to $\hat{s}\in{\sf Sol}(x)$. It implies that there is a bijection between ${\sf Sol}(y)$ and ${\sf Sol}(x)$.
The notion of fully-polynomial enumeration kernel(ization) is defined in a similar way. The only difference is that the condition (ii) of Definition 2.1 is replaced by the following condition.
(ii*)
For every $s\in{\sf Sol}(y)$, the solution-lifting algorithm ${\mathcal{A}}^{\prime}$ computes in time polynomial in $|x|+|y|+\kappa(x)+\kappa(y)$ a nonempty set of solutions $S_{s}\subseteq{\sf Sol}(x)$ such that $\{S_{s}\mid s\in{\sf Sol}(y)\}$ is a partition of ${\sf Sol}(x)$.
In the condition (ii), the solution-lifting algorithm for a fully-polynomial enumeration kernel runs in polynomial-time but the solution-lifting algorithm for a polynomial-delay enumeration kernelization runs in polynomial-delay. It is not very hard to observe that every bijective enumeration kernelization is a fully-polynomial enumeration kernelization and every fully-polynomial enumeration kernelization is a polynomial-delay enumeration kernelization [20]. Let $\Pi=(L,{\sf Sol},\kappa)$ be a parameterized enumeration problem such that $L$, the decision version of $\Pi$ is a decidable problem. Golovach et al. [20] have proved in their paper that (i) $\Pi$ admits an FPT-enumeration algorithm (an FPT-delay enumeration algorithm) if and only if $\Pi$ admits a fully-polynomial enumeration kernel (polynomial-delay enumeration kernel), and (ii) in fact, $\Pi$ can be solved in polynomial-time (or with polynomial-delay) if and only if $\Pi$ admits a fully-polynomial (polynomial-delay) enumeration kernelization of constant size.
Notice that the enumeration kernelizations that can be inferred from the above does not necessarily have polynomial size. For the decision versions of a parameterized problem, there is a well-known framework of cross composition [5] to refute the existence of polynomial kernels unless ${\sf NP\subseteq coNP/poly}$. No similar framework is yet known for parameterized enumeration problems. But, it is not very hard to see that if a parameterized enumeration problem $\Pi=(L,{\sf Sol},\kappa)$ admits a polynomial-delay enumeration kernel of polynomial size, then its decision version $L$ admits a polynomial kernel when parameterized by $\kappa$. Therefore, considering $L$ a decidable problem such that $\Pi=(L,{\sf Sol},\kappa)$ is a parameterized enumeration problem, if $L$ parameterized by $\kappa$ does not admit a polynomial kernel unless ${\sf NP\subseteq coNP/poly}$, then $\Pi$ does not admit a polynomial-delay/fully-polynomial enumeration kernel of polynomial size unless ${\sf NP\subseteq coNP/poly}$.
We would like to highlight that Gomes and Sau [15] have refuted the existence of polynomial kernel for Degree-$d$-Cut when parameterized solely by the treewidth (${\sf tw}$) or by cliquewidth of the input graph unless ${\sf NP\subseteq coNP/poly}$. Therefore, the following statement holds true.
Lemma 2.2.
For any fixed $d\geq 1$, Enum Deg-$d$-Cut, Enum Max Deg-$d$-Cut and Enum Min Deg-$d$-Cut do not admit any polynomial-delay enumeration kernelizations of polynomial size when parameterized by the treewidth (or cliquewidth respectively), unless ${\sf NP\subseteq coNP/poly}$.
3 Enumeration Kernelizations Parameterized by Vertex Cover Number
In this section, we consider the vertex cover number (vc) as the parameter and $d$ is a fixed positive integer. The size of a vertex cover with minimum cardinality is the vertex cover number of the input graph $G$ and is denoted by ${\textsf{vc}}(G)$. We can exclude this assumption that an explicit vertex cover is given with the input. We simply invoke a polynomial-time algorithm [31] that outputs a vertex cover $S$ with at most $2{\textsf{vc}}(G)$ vertices. Hence, we consider $(G,S,k)$ as the input instance such that $S$ is a vertex cover of $G$ such that $|S|=k\leq 2{\textsf{vc}}(G)$. The input instance for both the other variants Enum Min Deg-$d$-Cut and Enum Max Deg-$d$-Cut remains identical. Suppose that $I=V(G)\setminus S$. Let $I_{d}$ be the vertices of $I$ that has degree at most $d$ in $G$, and $I_{d+1}$ be the vertices of $I$ that has degree at least $d+1$ in $G$.
3.1 Marking Scheme and Kernelization Algorithm
We first describe a marking procedure ${\sf MarkVC}(G,S,k)$, that will be used by all the kernelizations in this section.
Figure 2: An illustration of attaching pendant cliques for $d=2$ with the set $T=\{x,y\}\subseteq S$. The green, blue, and pink colored vertices are the false twins of $u$ such that $N_{G}(u)=T$. The marked false twins are the red and the blue vertices. The pendant clique $C_{u}$ contains 6 vertices including $u$ and $C_{1}$ contains 7 vertices including $u_{1}$.
Procedure: ${\sf MarkVC}(G,S,k)$.
  1. 1.
    Initialize $H^{*}:=G$ and $S_{0}:=\emptyset$.
  2. 2.
    Mark one arbitrary isolated vertex of $H^{*}$ if exists.
  3. 3.
    For every nonempty set $T\in{{S}\choose{\leq d}}$, let there exists $u\in I_{d}$ such that $N_{G}(u)=T$. If $u$ has at most $2d$ false twins in $I_{d}$, then mark $u$ and all its false twins from $I_{d}$.
  4. 4.
    For every nonempty set $T\in{{S}\choose{\leq d}}$, let there is $u\in I_{d}$ such that $N_{G}(u)=T$. If $u$ has more than $2d$ false twins in $I_{d}$, then mark $u$ and $12d+$ 1 false twins of $u$ from $I_{d}$. Let $u_{1},\ldots,u_{d-1}\in I_{d}$ be $d-1$ arbitrary marked false twins of $u$. If $u$ has an unmarked false twin, then add a pendant clique $C_{u}$ with $22d+$ 2 vertices into $u$ and set $S_{0}:=S_{0}\cup(C_{u}\setminus\{u\})$. Then, for every $i\in[d-1]$, add a pendant clique $C_{i}$ with $22d+i+$ 2 vertices into $u_{i}$ and set $S_{0}:=S_{0}\cup(C_{i}\setminus\{u_{i}\})$. See Figure 2 for an illustration with $d=2$ and $|T|=d$.
  5. 5.
    For every nonempty set $T\in{{S}\choose{d+1}}$ mark $\min\{2d+1,|\bigcap\limits_{x\in T}N_{G}(x)\cap I_{d+1}|\}$ arbitrary vertices from $\bigcap\limits_{x\in T}N_{G}(x)\cap I_{d+1}$. Informally, for every $T\in{{S}\choose{d+1}}$, mark $12d+$ 1 vertices from $I_{d+1}$ that are common neighbors of the vertices of $T$.
We use $Z$ to denote the set of all marked vertices from $I$ and let $H=H^{*}[S\cup Z\cup S_{0}]$. It is straightforward to see that $H$ can be constructed in polynomial-time. The following observation illustrates some important properties of $H$.
Observation 3.1.
The number of vertices in $S_{0}$ is $\mathcal{O}(d^{3}|S|^{d})$ and every connected component of $H[S_{0}]$ is a clique with at least $12d+$ 1 and at most $3d$ vertices. Moreover, $S\cup S_{0}$ is a vertex cover of $H$ and the total number of vertices in $H$ is at most $18d^{3}|S|^{d+1}$.
Proof.
Observe that $|Z|\leq 1+(2d+2)\sum\limits_{i=1}^{d+1}{{|S|}\choose{i}}$ implying $Z$ has $\mathcal{O}(d|S|^{d+1})$ vertices. In particular $|Z|\leq 6d^{2}|S|^{d+1}$. Moreover, for every $T\in{{S}\choose{d}}$, the construction ensures that there are at most $d$ vertices from $I_{d}$ whose neighborhood in $G$ is exactly $T$ and there is a pendant clique with at at least $22d+$ 2 vertices and most $13d+$ 1 vertices in $S_{0}$ attached to it. In particular, every attached pendant clique $C$ contains one vertex from $I_{d}$ and the other vertices in $S_{0}$. Hence, every connected component of $S_{0}$ has at least $12d+$ 1 vertices and at most $3d$ vertices. Therefore, the total number of vertices in $H$ is at most $18d^{3}|S|^{d+1}$ and the number of vertices in $S_{0}$ is $\mathcal{O}(d^{3}|S|^{d})$. Also, as $S$ is a vertex cover of $G$ and by construction, $Z\subseteq I$, it follows that $S\cup S_{0}$ is a vertex cover of $H$. ∎
For the rest of the section, we let $\widetilde{H}=H-S_{0}$. In the case of enumeration kernelization, it is not sufficient to prove that the output instance is equivalent to the input instance. We need to prove some additional properties that are essential to establish the information of all $d$-cuts of $G$ are stored in the output graph $H$ which is the kernel. First we prove that any $d$-cut of $H$ and is a $d$-cut of $\widetilde{H}$ and vice versa. Then, we prove some structural properties that establish the relation between the $d$-cuts of $\widetilde{H}$ and the $d$-cuts of $G$. We use the Lemma 2.1 to prove the following observation.
Observation 3.2.
Let $F$ be a $d$-cut of $H$ and $uv\in E(H)$ such that $u,v\in S_{0}$ or $u\in S_{0},v\in Z$, then $uv\notin F$.
Proof.
Let the premise of the statement be true and $uv\in E(H)$ such that either $u,v\in S_{0}$ or $u\in S_{0},v\in Z$. It follows from the Observation 3.1 that every connected component of $H[S_{0}]$ is a clique with at least $12d+$ 1 vertices and at most $3d$ vertices. Furthermore, if $u,v\in S_{0}$ then $u,v\in C$ for a connected component $C$ of $H[S_{0}]$ that is a clique at least $12d+$ 1 vertices. Additionally, if $u\in S_{0}$ and $v\in Z$, then $u,v\in\widehat{C}$ for a pendant clique with at least $22d+$ 2 vertices attached to $v$. Let $(A,B)$ be the corresponding $d$-cut of $H$ such that $F=E_{H}(A,B)$. Since $(A,B)$ is a $d$-cut of $H$ and both $uv$ is an edge of a clique $\widehat{C}$ containing at least $12d+$ 1 vertices, it follows from Lemma 2.1 that either $\widehat{C}\subseteq A$ or $\widehat{C}\subseteq B$. Therefore, $uv\notin F$. ∎
The above observation implies that if an edge of $H$ is incident to the vertices of $S_{0}$, then that edge cannot be present in any $d$-cut of $H$. We use Lemma 2.1, Observations 3.1, and 3.2 to prove the following statement, saying that the collection of all $d$-cuts of $H$ and the collection of all $d$-cuts of $\widetilde{H}$ are precisely the same.
Lemma 3.1.
A set of edges $F$ is a $d$-cut of $\widetilde{H}$ if and only if $F$ is a $d$-cut of $H$.
Proof.
We first give the backward direction ($\Leftarrow$) of the proof. Let $(A,B)$ denotes the bipartition such that $F=E_{H}(A,B)$ is a $d$-cut of $H$. Due to Observation 3.1, any connected component $C$ of $H[S_{0}]$ is a clique containing at least $12d+$ 1 vertices. Hence, it follows from Lemma 2.1 that either $C\subseteq A$ or $C\subseteq B$. Furthermore, it follows from Observation 3.2 that if an edge $uv\in E(H)$ is incident to the vertices of $S_{0}$, then $uv\notin F$. Hence, the edges $E_{H}(A\setminus S_{0},B\setminus S_{0})$ and the edges $E_{H}(A,B)$ are precisely the same, implying that $F$ is a $d$-cut of $\widetilde{H}$.
For the forward direction ($\Rightarrow$), let $(A,B)$ be the bipartition corresponding to the $d$-cut $F=E_{\widetilde{H}}(A,B)$ of $\widetilde{H}$. Note that every connected component of $H[S_{0}]$ is a clique with at least $12d+$ 1 vertices. Furthermore, if $C$ is a connected component of $H[S_{0}]$, then every vertex of $C$ is adjacent to a unique vertex $u\in I_{d}$. Therefore, if $u\in A$, then we add the vertices of $C$ into $A$. Otherwise, $u\in B$, then we add the vertices of $C$ into $B$. This completes our construction of a $d$-cut of $H$ and the set of edges remain the same as $E_{\widetilde{H}}(A,B)$. This completes the proof that $F$ is also a $d$-cut of $H$. ∎
From the above lemma, it follows that that any minimal/maximal $d$-cut of $H$ is a minimal/maximal $d$-cut of $\widetilde{H}$ and vice versa. Therefore, it is sufficient to focus on establishing relationship between the $d$-cuts of $\widetilde{H}$ and the $d$-cuts of $G$.
3.2 Structural Properties of $d$-cuts
In order to establish relations of the $d$-cuts of $G$ and $\widetilde{H}$, we have to prove some important structural properties of the $d$-cuts of $\widetilde{H}$ and the $d$-cuts of $G$. For every nonempty $T\in{{S}\choose{\leq d}}$, we first define $L_{T}=\{xy\in E(G)\mid x\in T,y\in I_{d},N_{G}(y)=T\}$ and let $L=\cup_{T\in{{S}\choose{\leq d}}}L_{T}$. Informally, $L_{T}$ is the set of all edges that have one endpoint in $T$ and other endpoint in $I_{d}$ such that if $xy\in L_{T}$ and $y\in I_{d}$, then $x\in T$ and $N_{G}(y)=T$. We call a nonempty set $T\in{{S}\choose{\leq d}}$ good if there are at most $12d+$ 1 vertices in $I_{d}$ whose neighborhood exactly equals $T$. Precisely, if |$|\{u\in I_{d}\mid N_{G}(u)=T\}|\leq 2d+1$, then $T$ is a good subset of $S$. A nonempty set $T\in{{S}\choose{\leq d}}$ is bad if $T$ is not a good subset of $S$. Observe from the MarkVC($G,S,k$) that if a nonempty set $T\in{{S}\choose{\leq d}}$ is good in $G$, then for every $u\in I_{d}\cap V(G)$ such that $N_{G}(u)=T$, every false twin of $u$ is marked in $Z$. Therefore, $u$ and each of its false twins are present in $\widetilde{H}$. Hence, such a subset $T\subseteq S$ is a good subset in $\widetilde{H}$ as well. Similarly, if a nonempty set $T\in{{S}\choose{\leq d}}$ is bad, then exactly $22d+$ 2 vertices from $I_{d}\cap V(G)$ each of which have neighborhood exactly $T$ are marked and are present in $\widetilde{H}$. Then, $T$ is a bad subset of $S$ in $\widetilde{H}$ as well. Let ${\mathcal{Y}}\subseteq{{S}\choose{d}}$ be the collection of all bad subsets of $S$ in $G$ (and in $\widetilde{H}$). We define $L({\mathcal{Y}})=\cup_{T\in{\mathcal{Y}}}L_{T}$ and $I({\mathcal{Y}})$ the set of all vertices of $I_{d}$ each of which have neighborhood exactly $T$ such that $T\in{\mathcal{Y}}$. Formally, $I({\mathcal{Y}})=\{u\in I_{d}\mid T=N_{G}(u)$ and $T\in{\mathcal{Y}}\}$. We will use Lemma 2.1 to prove that the marking scheme preserves all the $d$-cuts of $G-I({\mathcal{Y}})$.
Lemma 3.2.
A set of edges $F\subseteq E(G)$ is a $d$-cut of $G-I({\mathcal{Y}})$ if and only if $F$ is a $d$-cut of $\widetilde{H}-I({\mathcal{Y}})$.
Proof.
We can assume without loss of generality that there is $v\in V(G)\setminus(V(\widetilde{H})\cup I({\mathcal{Y}}))$ that is not marked by the procedure ${\sf MarkVC}(G,S,k)$. Otherwise, all the vertices of $I_{d+1}\cap V(G)$ are marked by our marking procedure ${\sf MarkVC}(G,S,k)$ implying that $V(G-I({\mathcal{Y}}))=V(\widetilde{H}-I({\mathcal{Y}}))$. Then, $G-I({\mathcal{Y}})$ and $\widetilde{H}-I({\mathcal{Y}})$ are the same graphs, and the lemma statement trivially holds true. So we include the above assumption (i.e., there is some unmarked vertex in $I_{d+1}\cup I({\mathcal{Y}})$) for both directions of the proof. For simplicity of presentation, we let $\vec{H}=\widetilde{H}-I({\mathcal{Y}})$ and $\vec{G}=G-I({\mathcal{Y}})$.
First, we give the backward direction ($\Leftarrow$) of the proof. Suppose that $(A^{\prime},B^{\prime})$ be the bipartition corresponding to a $d$-cut $F$ of $\vec{H}$. Consider a vertex $v$ that is in $\vec{G}$ but not in $\vec{H}$. Then, such a vertex $v$ satisfies that $v\in I_{d+1}\cap(V(G)\setminus V(\widetilde{H}))$. Hence, $v$ is a vertex of $I$ with at least $d+1$ neighbors in $S$ such that $v\notin V(\widetilde{H})$. Then, for every $T\subseteq N_{G}(v)$ with $|T|=d+1$ the procedure ${\sf MarkVC}(G,S,k)$ did not mark $v$ from the vertices of $\bigcap\limits_{x\in T}(N_{G}(x)\cap I_{d+1})$. What it means is that $\widetilde{H}$ has $12d+$ 1 vertices in $I_{d+1}$, the neighborhood of each of which contains $T$. If there are two distinct subsets $T_{1},T_{2}\subseteq N_{G}(v)$ with $d+1$ vertices each, then the procedure ${\sf MarkVC}(G,S,k)$ did not mark the vertex $v$ from the vertices of $\bigcap\limits_{x\in T_{1}}(N_{G}(x)\cap I_{d+1})$ and $\bigcap\limits_{x\in T_{2}}(N_{G}(x)\cap I_{d+1})$. The above implies that $H$ has $12d+$ 1 (marked) vertices, the neighborhood of each of which contains $T_{1}$ as well as $T_{2}$. Then by Lemma 2.1, either $T_{1}\subseteq A^{\prime}$ or $T_{1}\subseteq B^{\prime}$. Moreover, it also holds from Lemma 2.1 that either $T_{2}\subseteq A^{\prime}$ or $T_{2}\subseteq B^{\prime}$. We have to argue that $T_{1}\cup T_{2}\subseteq A^{\prime}$ or $T_{1}\cup T_{2}\subseteq B^{\prime}$. For the sake of contradiction, let $T_{1}\subseteq A^{\prime}$ but $T_{2}\subseteq B^{\prime}$. In such a case, we construct a set $T_{3}$ by taking $d$ vertices from $T_{1}$ and one vertex of $T_{2}\setminus T_{1}$. Observe that $T_{3}\subseteq N_{G}(v)$ and has $d+1$ vertices. Then this assumption on $T_{3}$ implies that $T_{3}\cap A^{\prime},T_{3}\cap B^{\prime}\neq\emptyset$.
Note that $v\in\bigcap\limits_{x\in T_{3}}(N_{G}(x)\cap I_{d+1})$ and the procedure ${\sf MarkVC}(G,S,k)$ did not mark $v$ from the vertices of $\bigcap\limits_{x\in T_{3}}(N_{G}(x)\cap I_{d+1})$. It means that there are at least $12d+$ 1 vertices in $I$ the neighborhood of which contains $T_{3}$. But then from Lemma 2.1, it follows that $T_{3}\subseteq A^{\prime}$ or $T_{3}\subseteq B^{\prime}$. This contradicts the choice of $T_{3}$ satisfying $T_{3}\cap A^{\prime},T_{3}\cap B^{\prime}\neq\emptyset$. Therefore, all subsets of ${{N_{G}(v)}}$ with $d+1$ vertices either are contained in $A^{\prime}$ or is contained in $B^{\prime}$ implying that $N_{G}(v)\subseteq A^{\prime}$ or $N_{G}(v)\subseteq B^{\prime}$.
We provide a construction of a cut $(A,B)$ of $\vec{G}$ such that $F=E_{\vec{G}}(A,B)$. At the very beginning, we initialize $A:=A^{\prime}$ and $B:=B^{\prime}$. If $N_{G}(v)\subseteq A^{\prime}$, then we add $v$ into $A$, otherwise $N_{G}(v)\subseteq B^{\prime}$, and we add $v$ into $B$. This construction of $A$ and $B$ preserves that $E_{\vec{G}}(A,B)=E_{\vec{H}}(A^{\prime},B^{\prime})=F$. Therefore, $(A,B)$ is a $d$-cut of $\vec{G}$ such that $F=E_{\vec{G}}(A,B)$. This completes the correctness proof of the backward direction.
Let us now give the forward direction ($\Rightarrow$) of the proof. Suppose that $(A,B)$ be the corresponding vertex bipartition of a $d$-cut of $\vec{G}$ such that $F=E_{\vec{G}}(A,B)$. We argue that $F$ is a $d$-cut of $\vec{H}$ as well. For the sake of contradiction, we assume that there is $uv\in E(\vec{G})\setminus E(\vec{H})$ such that $uv\in F$. Since $\vec{H}$ is an induced subgraph of $\vec{G}$, then either $u\notin V(\vec{H})$ or $v\notin V(\vec{H})$. Without loss of generality, we assume that $u\notin V(\vec{H})$. Then, $u$ is one of the unmarked vertices from $I_{d+1}$, and we assume that $u\in A$. Consider an arbitrary $T\in{{N_{G}(u)}\choose{d+1}}$ such that $v\in T$. Corresponding to the set $T$, there are at least $12d+$ 1 marked vertices in $V(G)\cap I_{d+1}$ and $u$ is excluded from marking by the procedure ${\sf MarkVC}(G,S,k)$. In particular, there are at least $22d+$ 2 vertices in $G$ the neighborhood of which contains $T$. Then, due to Lemma 2.1, it follows that either $T\subseteq A$ or $T\subseteq B$. But, by our choice $u\in A$. Since $v\in T$ and $uv\in F$, it must be that $v\in B$. But then $T\subseteq B$ implies that $u\in A$ has at least $d+1$ neighbors in $B$. This contradicts to our initial assumption that $(A,B)$ is a $d$-cut of $\vec{G}$. This completes the proof that $F$ is also a $d$-cut of $\vec{H}$. ∎
Sometimes we say “$T$ is a bad set” to mean that “$T$ is a bad subset of $S$” (in the context of the graph $G$ or $\widetilde{H}$). Given an arbitrary subset $T\in{{S}\choose{\leq d}}$ and $u\in I_{d}$ such that $N_{G}(u)=T$, we define $J_{T}^{u}=\{ux\mid x\in T,u\in I_{d},N_{G}(u)=T\}$. That is, the set $J^{u}_{T}$ is the set of edges with one endpoint $u$ and the other endpoint in $T$ such that $N_{G}(u)=T$. A bad subset $T\in{{S}\choose{\leq d}}$ is called broken by an edge set $F$ if there exists $u\in I_{d}$ satisfying $N_{G}(u)=T$ such that $J^{u}_{T}\not\subset F$ and $J^{u}_{T}\cap F\neq\emptyset$. A bad subset $T\in{{S}\choose{\leq d}}$ is called occupied by an edge set $F\subseteq E(G)$ if $T$ is not broken by $F$. Equivalently, if a bad subset $T$ is occupied by an edge set $F$, then either $J^{u}_{T}\subseteq F$, or $J^{u}_{T}$ is disjoint from $F$. The following observation about bad subsets of $S$ can be proved by using Lemma 2.1.
Observation 3.3.
Let an edge set $F$ be a $d$-cut of $G$. Then, every bad subset $T\in{{S}\choose{\leq d}}$ is occupied by $F$.
Proof.
We give the proof for $G$. Similar arguments can be used to prove the analogous statement in $\widetilde{H}$. Let $(A,B)$ be the vertex bipartition of a $d$-cut $F$ of $G$ such that $F=E_{G}(A,B)$ and $T\in{{S}\choose{\leq d}}$ be a bad subset of $S$. If $|T|=1$, then clearly $T\subseteq A$ (or $T\subseteq B$). Hence, for any $u\in I_{d}$ with $N_{G}(u)=T$ satisfies that $J^{u}_{T}$ is either disjoint from $F$ or is contained in $F$. So, $T$ is occupied by $F$. When $|T|\geq 2$, it follows from Lemma 2.1 that $T\subseteq A$ (or $T\subseteq B$). We assume without loss of generality that $T\subseteq A$. Then, for any $u\in I_{d}$ satisfying $N_{G}(u)=T$, if $u\in A$, then $J^{u}_{T}\cap F=\emptyset$. If $u\in B$, then $J^{u}_{T}\subseteq F$. This completes the proof that $T$ is occupied by $F$. ∎
As a remark, we would like to mention that the converse of Observation 3.3 does not necessarily hold true. With the above observation in hand, we define the notion of equivalence between two edge sets as follows. Two sets of edges $F_{1}$ and $F_{2}$ are equivalent if the following properties are satisfied.
  1. (i)
    $F_{1}\setminus L=F_{2}\setminus L$,
  2. (ii)
    for every $T\in{{S}\choose{\leq d}}$, $|F_{1}\cap L_{T}|=|F_{2}\cap L_{T}|$,
  3. (iii)
    for every good subset $T\in{{S}\choose{\leq d}}$, $F_{1}\cap L_{T}=F_{2}\cap L_{T}$.
  4. (iv)
    a bad subset $T\in{{S}\choose{\leq d}}$ is occupied by $F_{1}$ if and only if $T$ is occupied by $F_{2}$.
It is clear from the definition that the above relation between two different edge sets is an equivalence relation. This definition not only holds for two edge sets of $G$, but also holds for two edge sets of $\widetilde{H}$ as well as for one edge set of $G$ and one edge set of $\widetilde{H}$. Since the edges of $L$ only involve the edges that have one endpoint in $S$ and one endpoint in $I$, the edges $F\setminus L$ include the edges of $F$ that have both endpoints in $S$. It is intuitive that in case two edge sets $F_{1}$ and $F_{2}$ are equivalent to each other, then both have the same intersection with the edges that have both endpoints in $S$. We use Lemma 2.1 and Observation 3.3 to prove the following lemma that illustrates some structural properties of equivalent sets of edges.
Lemma 3.3.
Let $F^{\prime}\subseteq E(\widetilde{H})$ be a $d$-cut (minimal or maximal $d$-cut, respectively) of $\widetilde{H}$. Then, every $F\subseteq E(G)$ that is equivalent to $F^{\prime}$ is a $d$-cut (minimal or maximal $d$-cut, respectively) of $G$.
Proof.
Let $(A^{\prime},B^{\prime})$ be the vertex bipartition corresponding to the $d$-cut $F^{\prime}$ of $\widetilde{H}$ such that $F^{\prime}=E(A^{\prime},B^{\prime})$ and let $F\subseteq E(G)$ be equivalent to $F^{\prime}$. It suffices to prove that for every $u\in V(G)$, there are at most $d$ edges of $F$ that are incident to $u$. The following three cases arise for a vertex $u\in V(G)$.
Case (i):
Let $u\in S$. Observe that the edges of $F$ that are incident to $u$ can contribute to $F\cap L_{T}$ for a nonempty set $T\in{{S}\choose{\leq d}}$, or to the set $F\setminus L$.
For a nonempty set $T\subseteq S$ and $u\in T$, we define $F_{T,u}=\{uv\in F\mid v\in I,N_{G}(v)=T\}$ and $F^{\prime}_{T,u}=\{uv\in F^{\prime}\mid v\in I,N_{\widetilde{H}}(v)=T\}$. Finally, define $F_{S,u}=\{uv\in F\mid v\in S\}$ and $F^{\prime}_{S,u}=\{uv\in F^{\prime}\mid v\in S\}$.
We argue by these sets one by one. Since $F$ is equivalent to $F^{\prime}$, it follows from the definition that that $F^{\prime}\cap E(S)$ contribute to the edges of $F\setminus L_{T}$. Since $F$ is equivalent to $F^{\prime}$, it follows that $F^{\prime}\cap E(S)=F\cap E(S)$. It implies that $F^{\prime}_{S,u}=F_{S,u}$. Similarly, if $T\in{{S}\choose{\leq d}}$ is a nonempty good subset of $S$, then $F\cap L_{T}=F^{\prime}\cap L_{T}$. Therefore, it must be that $F_{T,u}=F^{\prime}_{T,u}$. If $T\in{{S}\choose{\geq d+1}}$, then notice that $F_{T,u}\subseteq F\setminus L$ and $F^{\prime}_{T,u}\subseteq F^{\prime}\setminus L$. Since $F\setminus L=F^{\prime}\setminus L$ due to the equivalence between $F$ and $F^{\prime}$, it must hold that $F_{T,u}=F^{\prime}_{T,u}$.
Finally, we consider when $T\in{{S}\choose{\leq d}}$ is a nonempty bad subset of $S$. It follows from the Lemma 2.1 that $T\subseteq A^{\prime}$ or $T\subseteq B^{\prime}$. Without loss of generality, we assume that $T\subseteq A^{\prime}$. Then let $I_{T}\subseteq I_{d}\cap B^{\prime}$ be the vertices whose neighborhood is exactly the same as $T$ and let $\alpha_{T}=|I_{T}|$. Then, we look at the edges of $F_{T,u}$ and $F^{\prime}_{T,u}$. Note that $|I_{T}|=|F^{\prime}_{T,u}|=\alpha_{T}$ since $u\in T\subseteq A^{\prime}$. Since $F$ is equivalent to $F^{\prime}$, it follows that $|F\cap L_{T}|=|F^{\prime}\cap L_{T}|$. Moreover, as $T$ is a bad subset of $S$ with at most $d$ vertices and $F^{\prime}$ is a $d$-cut of $\widetilde{H}$, it follows from Observation 3.3 that $T$ is occupied by $F^{\prime}$. As $F$ is equivalent to $F^{\prime}$, it follows that $T$ is occupied by $F$ as well. Hence, for any $w\in I_{d}$ satisfying $N_{G}(w)=T$, it must be that either $J_{T}^{w}\subseteq F$ or $J_{T}^{w}\cap F=\emptyset$. Since $|F^{\prime}\cap L_{T}|=|F\cap L_{T}|$, there are exactly $\alpha_{T}$ edges incident to $u$ that are in $F_{T,u}$. Hence, $|F_{T,u}|=\alpha_{T}$ implying that $|F_{T,u}|=|F^{\prime}_{T,u}|$.
Observe from the case analysis above that the number of edges in $F^{\prime}$ that are incident to $u$ is equal to the number of edges in $F$ that are incident to $u$. Hence, there are at most $d$ edges in $F$ that are incident to $u$.
Case (ii):
Let $u\in I$. There are two subcases, $u\in I_{d}$ and $u\in I_{d+1}$. If $u\in I_{d}$ and $u$ has no neighbor in $S$, then $u$ is isolated in $G$ (and in $\widetilde{H}$), hence there cannot exist any edge in $F$ incident to $u$. As $F$ is equivalent to $F^{\prime}$, there cannot exist any edge in $F^{\prime}$ that is incident to $u$.
Suppose that $u$ has some neighbor in $S$ and let $T=N_{G}(u)$. Note that $T\neq\emptyset$ and has at most $d$ vertices. If $T$ is a good subset of $S$ (in $G$ and in $\widetilde{H}$), then $u\in V(\widetilde{H})$. Since $F$ is equivalent to $F^{\prime}$, it holds that $F\cap L_{T}=F^{\prime}\cap L_{T}$. Hence, $F_{T,u}=F^{\prime}_{T,u}$. On the other hand, if $T$ is a bad subset of $S$, then $T$ is occupied by $F^{\prime}$. Since $F$ is equivalent to $F^{\prime}$, $T$ is occupied by $F$ as well. Hence, the set $J^{u}_{T}=\{uv\in E(G)\mid N_{G}(u)=T\}$ is either disjoint from $F$ or prpoerly contained in $T$. Since $u\in I_{d}$ and $N_{G}(u)=T$ and $|T|\leq d$, it follows that there are at most $d$ edges in $F$ that are incident to $u$.
The last subcase is when $u\in I_{d+1}$. There are two subcases. Either $u\in V(\widetilde{H})$ or $u\in V(G)\setminus V(\widetilde{H})$. Note that the edges of $F^{\prime}$ that are incident to $u$ contribute to the set $F^{\prime}\setminus L$. Since $F$ is equivalent to $F^{\prime}$, it holds that $F\setminus L=F^{\prime}\setminus L$. Hence, the set of edges in $F^{\prime}$ that are incident to $u$ is the same as the set of edges in $F$ that are incident to $u$. Since $F^{\prime}$ is a $d$-cut, there are at most $d$ edges of $F^{\prime}$. Therefore, the same set of at most $d$ edges of $F$ are incident to $u$.
Since the above mentioned cases are mutually exhaustive, this completes the proof of the lemma. ∎
The intuition of above lemma is that given a $d$-cut $F^{\prime}$ of $\widetilde{H}$, the solution-lifting algorithm should consider to enumerate the collection of edge sets of $G$ that are equivalent to $F^{\prime}$. But the following lemma is a stronger statement the proof of which uses the Lemma 2.1, Lemma 3.2 and Lemma 3.3.
Lemma 3.4.
An edge set $F\subseteq E(G)$ is a $d$-cut (minimal or maximal $d$-cut respectively) of $G$ if and only if $\widetilde{H}$ has a $d$-cut (minimal or maximal $d$-cut respectively) $F^{\prime}$ that is equivalent to $F$.
Proof.
Let us first give the backward direction ($\Leftarrow$) of the proof. Consider a $d$-cut $F^{\prime}$ of $\widetilde{H}$ and let $F=F^{\prime}$. It follows that $F$ is equivalent to $F^{\prime}$. Then, by Lemma 3.3, $F$ is a $d$-cut of $G$ as well.
For the forward direction ($\Rightarrow$), let $(A,B)$ be a $d$-cut of $G$ such that $F=E_{G}(A,B)$. If $F=\emptyset$, then $G$ is disconnected. Due to the procedure ${\sf MarkVC}(G,S,k)$, it follows that $\widetilde{H}$ is also disconnected, and the set of edges $F^{\prime}=F=\emptyset$ is a $d$-cut of $\widetilde{H}$ as well. Then $F^{\prime}$ is equivalent to $F$, and the statement holds. Otherwise $F\neq\emptyset$.
We construct $F^{\prime}$ with vertex bipartition $(A^{\prime},B^{\prime})$ as follows. Initialize $A^{\prime}:=A$ and $B^{\prime}:=B$. We consider $T\in{{S}\choose{\leq d}}$ such that $F\cap L_{T}\neq\emptyset$. Let ${\mathcal{Y}}$ be the collection of subsets of $S$ that are bad and $I({\mathcal{Y}})$ denotes the vertices of $I_{d}$ the neighborhood of which is $T$ for some $T\in{\mathcal{Y}}$. If $F\cap L_{T}=\emptyset$ for all $T\in{\mathcal{Y}}$, then $F\subseteq E(G-I({\mathcal{Y}}))$. Then, $F$ is a $d$-cut of $G-I({\mathcal{Y}})$. It follows from Lemma 3.2 that $F$ is a $d$-cut of $\widetilde{H}-I({\mathcal{Y}})$. Furthremore, from Observation 3.3, every $T\in{\mathcal{Y}}$ is occupied by $F$. As $T$ is occupied by $F$ and $F\cap L_{T}=\emptyset$, it implies for every $u\in I({\mathcal{Y}})$ with $N_{G}(u)=T$ that $T\subseteq A$ if and only if $u\in A$. Then, setting $A^{\prime}=A\setminus I({\mathcal{Y}})$ and $B^{\prime}=B\setminus I({\mathcal{Y}})$ ensures us that the set of edges crossing the cut $(A^{\prime},B^{\prime})$ remains the same as $F$. Hence, $F^{\prime}=F$ is a $d$-cut of $\widetilde{H}$.
Otherwise, there is a bad subset $T$ such that $F\cap L_{T}\neq\emptyset$. Then, $I$ has more than $12d+$ 1 vertices in $G$, each of which has neighborhood exactly the same as $T$. If $T=\{x\}$, then there are $12d+$ 1 pendant neighbors of $x$ apart from $u$ that are marked by ${\sf MarkVC}(G,S,k)$. Let $x\in A$ and suppose that there are at most $d^{*}\leq d$ pendant neighbors of $x$ that are in $B$. Replace these $d^{*}$ neighbors of $x$ from $B$ with some other $d^{*}$ marked pendant neighbors from $I_{d}$ in $B^{\prime}$. Hence, for a bad set $T$ with only one vertex, it follows that $|F\cap L_{T}|=|F^{\prime}\cap L_{T}|$. Additionally, it also preserves that $\{x\}$ is occupied by $F^{\prime}$. Otherwise, $2\leq|T|\leq d$. Since $T$ is a bad subset, there are at least $22d+$ 2 vertices from $I_{d}$ each of which has neighborhood exactly same as $T$. From Lemma 2.1, either $T\subseteq A$ or $T\subseteq B$. If $T\subseteq A$, then there are at most $d^{*}\leq d$ vertices in $B$ each of which have neighborhood exactly $T$. We replace these $d^{*}$ vertices from $B$ by some other $d^{*}$ marked vertices from $Z$ into $B^{\prime}$. Each of these $d^{*}$ marked vertices that are put into $B^{\prime}$ also have the neighborhood exactly $T$ (in $\widetilde{H}$). Similarly, if $T\subseteq B$, then there are at most $d^{*}\leq d$ vertices in $A$. Same as before, we replace these $d^{*}$ vertices of $A^{\prime}$ with some $d^{*}$ marked vertices from $Z$ and put them into $A^{\prime}$. Each of these $d^{*}$ vertices have neighborhood exactly $T$.
We repeat this above mentioned procedure for all nonempty set $T\in{{S}\choose{\leq d}}$ that is bad. Once this replacement procedure is done, observe that the set of edges crossing $(A^{\prime},B^{\prime})$ are in $\widetilde{H}$. After that, from $A^{\prime}$ as well as from $B^{\prime}$, we delete all vertices of $I({\mathcal{Y}})$ that are in $V(G)\setminus V(\widetilde{H})$ and let $F^{\prime}=E_{\widetilde{H}}(A^{\prime},B^{\prime})$. Observe by construction that every replacement procedure ensures the following. If $T$ was occupied by $F$, then $T$ continues to remain occupied by $F^{\prime}$. Also, every replacement step preserves that $|F\cap L_{T}|=|F^{\prime}\cap L_{T}|$. Furthermore, for good subsets $T\in{{S}\choose{\leq d}}$, this procedure does not made any change to the edge set $F\cap L_{T}$. The replacement procedure also has not made any change to the edges of $F\setminus L$. Hence, $F^{\prime}$ is equivalent to $F$.
It remains to prove that $F^{\prime}$ is also a $d$-cut of $\widetilde{H}$. Let $F_{1}=F^{\prime}\cap L$ and $F_{2}=F^{\prime}\setminus L$. Either $F_{2}=\emptyset$ or $F_{2}\neq\emptyset$. If $F_{2}=\emptyset$, then $F^{\prime}=F_{1}$. Every step of replacement involves a bad subset $T\subseteq A$ or a $T\subseteq B$. If $T\subseteq A$ and there are $d^{*}$ vertices including some unmarked vertex in $B$, each of which having neighborhood exactly $T$, then replacement step replaces these $d^{*}(\leq d)$ vertices of $B$ by some other $d^{*}$ marked vertices into $B^{\prime}$. In particular, each of those $d^{*}$ marked vertices have neighborhood exactly equal to $T$. Moreover, two different replacement steps involve pairwise disjoint set of edges. Hence, in $F^{\prime}=(A^{\prime},B^{\prime})$, it preserves the property that for all $x\in A^{\prime}$, at most $d$ neighbors are in $B^{\prime}$, and vice versa. Hence, $F^{\prime}=F_{1}$ is a $d$-cut of $\widetilde{H}$.
For minimal and maximal $d$-cuts, the arguments are exactly the same. Let there are two $d$-cuts $F$ and $\hat{F}$ such that $F\subset\hat{F}$. The construction of the forward direction ($\Rightarrow$) of the proof ensures that $\widetilde{H}$ has two $d$-cuts $F^{\prime}$ and $\hat{F}^{\prime}$ that is equivalent to $F$ and $\hat{F}$ respectively satisfying the same inclusion property $F^{\prime}\subset\hat{F}^{\prime}$. This completes the proof of the lemma. ∎
3.3 Designing the Enumeration Kernelization Algorithms
Finally, we are ready to illustrate how we can use structural properties of equivalent $d$-cuts to design (polynomial-delay) enumeration kernelization algorithms. For minimal $d$-cuts, there are a few special properties that we need to understand. Let ${\mathcal{Y}}\subseteq{{S}\choose{\leq d}}$ be a collection of all bad subsets of $S$ in $G$. So, a $d$-cut $F\subseteq E(G)$ can be expressed as $F=F_{b}\uplus F_{g}$. In this partition, we assume that $F_{b}=\cup_{T\in{{\mathcal{Y}}}}(F\cap L_{T})$, and $F_{g}=F\setminus F_{b}$. We can also partition the edges $F^{\prime}\subseteq E(\widetilde{H})$ of a $d$-cut of $\widetilde{H}$ similarly. The following observation can be proved by Lemma 2.1 that illustrates a crucial property of any minimal $d$-cut of $G$ (or of $\widetilde{H}$).
Observation 3.4.
Let $F=(A,B)$ be a minimal $d$-cut of a connected graph $G$ (or of $\widetilde{H}$) and $F=F_{b}\uplus F_{g}$ be defined as before. If $F_{g}\neq\emptyset$, then $F_{b}=\emptyset$.
Proof.
We give the proof for the graph $G$, and the proof for the graph $\widetilde{H}$ will use similar arguments. Let $F=F_{b}\uplus F_{g}$ be a minimal $d$-cut such that $F_{g},F_{b}\neq\emptyset$ (for the sake of contradiction). Furthermore, we assume that the vertex bipartition for $F$ is $(A,B)$. Since $F_{b}\neq\emptyset$, there is $u\in I_{d}\cap A$ such that $N_{G}(u)\cap B\neq\emptyset$. But $|N_{G}(u)|\leq d$ and $N_{G}(u)$ is a bad subset of $S$. It follows from Lemma 2.1 that $N_{G}(u)\subseteq B$. But, $F_{g}\neq\emptyset$. Therefore, moving $u$ from $A$ to $B$ preserves the property that the resulting partition corresponds to a nonempty $d$-cut, the edge set of which is a proper subset of $F$. This keeps $F_{g}$ unchanged. Repeating this procedure for every vertex $u\in A\cap I_{d}$ (or $u\in B\cap I_{d}$) such that $N_{G}(u)\subseteq B$ (or $N_{G}(u)\subseteq A$) keeps $F_{g}$ unchanged while makes $F_{b}=\emptyset$. Then, we get a nonempty $d$-cut that is a proper subset of $F$. This contradicts the minimality of $F$ completing the proof of the observation. ∎
The above observation holds for $d$-cuts of $\widetilde{H}$ as well. In particular, if there is $u\in I_{d}$ such that $N_{\widetilde{H}}(u)$ is a bad subset and has at most $d$ vertices, then the cut ($(\{u\},V(H-(S_{0}\cup\{u\}))$ is a minimal $d$-cut of $\widetilde{H}$. The above observation implies that for every minimal $d$-cut of $\widetilde{H}$, either $F_{b}=\emptyset$ or $F_{g}=\emptyset$ or both. We use Lemma 3.3, Lemma 3.4, and Observation 3.4 to prove the main results of our paper.
See 1.1
Proof.
We prove the statements in the given order. For both the items, the enumeration kernelization has a kernelization algorithm and a solution-lifting algorithm.
  1. 1.
    Without loss of generality, we assume that $(G,S,k)$ be the input instance such that $|S|=k\leq 2{\textsf{vc}}(G)$ and $S$ is a vertex cover of $G$. We divide the situation into two cases.
    Case (i):
    If $G$ has at most $k^{d}$ vertices, the kernelization algorithm outputs $H=G$ itself. The solution-lifting algorithm takes a minimal $d$-cut $F^{\prime}$ of $H$ along with the input graph $G$ and output graph $H$, outputs $F=F^{\prime}$ only as the only minimal $d$-cut of $G$. Clearly both the two properties (i) and (ii) of fully-polynomial enumeration kernel are satisfied and the size of the kernel is $\mathcal{O}({\textsf{vc}}^{d})$.
    Case (ii):
    The other case is when $G$ has more than $k^{d}$ vertices. It means that $n>k^{d}$. The kernelization algorithm performs the marking scheme described by ${\sf MarkVC}(G,S,k)$ and let $H$ be the output graph. Output the instance $(H,S\cup S_{0},|S\cup S_{0}|)$. It follows from the Observation 3.1 that $S\cup S_{0}$ is a vertex cover of $H$ with $\mathcal{O}(d^{3}k^{d})$ vertices and $H$ has $\mathcal{O}(d^{3}k^{d+1})$ vertices. This completes the description of the kernelization algorithm.
    The solution-lifting algorithm works as follows. Let $F^{\prime}\subseteq E(H)$ be a minimal $d$-cut of $H$ and $\widetilde{H}=H-S_{0}$. It follows from Lemma 3.1 that any minimal $d$-cut $F^{\prime}$ of $H$ is a minimal $d$-cut of $\widetilde{H}$ and vice versa. Let ${\mathcal{Y}}\subseteq{{S}\choose{\leq d}}$ be the collection of all bad subsets of $S$ and $I({\mathcal{Y}})=\{u\in I_{d}\mid N_{G}(u)=T$ and $T\in{\mathcal{Y}}\}$. Suppose that $(A^{\prime},B^{\prime})$ is the vertex bipartition of $\widetilde{H}$ such that $F^{\prime}=E_{\widetilde{H}}(A^{\prime},B^{\prime})$. Let $F_{b}^{\prime}=\cup_{T\in{\mathcal{Y}}}(F^{\prime}\cap L_{T})$ and $F_{g}^{\prime}=F^{\prime}\setminus F_{b}$. If $F_{b}^{\prime}=\emptyset$, then $F^{\prime}=F_{g}^{\prime}$ is a minimal $d$-cut of $\widetilde{H}-I({\mathcal{Y}})$. By Lemma 3.2, $F^{\prime}$ is a $d$-cut of $G-I({\mathcal{Y}})$. As $F^{\prime}$ is equivalent to itself, by Lemma 3.3 and Lemma 3.4, $F^{\prime}$ is a minimal $d$-cut of $G$ as well. The solution-lifting algorithm just outputs $F=F^{\prime}$ as the only minimal $d$-cut of $G$.
    Otherwise, $F_{b}^{\prime}\neq\emptyset$. Then it follows from Observation 3.4 that $F_{g}^{\prime}=\emptyset$. Then, there is a bad subset $T$ of $S$ in $G$ such that $F^{\prime}\cap L_{T}\neq\emptyset$. Due to Lemma 2.1, $T\subseteq B^{\prime}$ or $T\subseteq A^{\prime}$. In addition, it also follows from Observation 3.3 that $T$ is occupied by $F^{\prime}$. Then, if there is $u\in I_{d}\cap A^{\prime}$ such that $J^{u}_{T}\subseteq F^{\prime}$. Therefore, by the property of minimality, it follows that if $T\subseteq B^{\prime}$, then $A^{\prime}=\{u\}$. Alternatively if $T\subseteq A^{\prime}$, then $B^{\prime}=\{u\}$. Hence, $F_{b}=F^{\prime}$. We describe the solution-lifting algorithm that $A=\{u\}$ and $B=V(\widetilde{H})\setminus\{u\}$. Let $C_{u}$ be the pendant clique attached $u$ in $H$ such that $C_{u}\setminus\{u\}\subseteq S_{0}$. Observe that $N_{G}(u)=N_{\widetilde{H}}(u)=T$. If $C_{u}$ is a clique with exactly $22d+$ 2 vertices, then the solution-lifting algorithm outputs the $d$-cuts $(\hat{A},\hat{B})$ such that $\hat{B}=B$ and $\hat{A}=\{v\}$ for all $v\in(I_{d}\setminus V(H))\cup\{u\}$ such that $v$ is a false twin of $u$. Otherwise, if $C_{u}$ is a pendant clique attached to $u$ containing more than $22d+$ 2 vertices or no pendant clique is attached to $C_{u}$, then the solution-lifting algorithm outputs only $F=F^{\prime}$, i.e. the same $d$-cut of $G$. Since $F$ is a minimal $d$-cut of $\widetilde{H}$ and $F$ is equivalent to $F^{\prime}(=F)$, due to Lemma 3.3, it is also minimal a $d$-cut of $G$. Also observe that $F$ is the only minimal $d$-cut of $G$ that is equivalent to $F^{\prime}$. Observe that these $d$-cuts are all equivalent to $F^{\prime}$ and can be outputted in polynomial-time. Moreover, for two distinct $d$-cuts of $H$ (and hence of $\widetilde{H}$), the collections of $d$-cuts of $G$ that are outputted are pairwise disjoint. Furthermore, due to Lemma 3.3, for every $d$-cut of $G$, there is an equivalent $d$-cut in $\widetilde{H}$ and by Lemma 3.1, that is also a $d$-cut of $H$. Since the solution-lifting algorithm runs in polynomial-time, this satisfies the condition (ii) of fully-polynomial enumeration kernel.
  2. 2.
    We give a proof of this statement for Enum Deg-$d$-Cut. The proof for Enum Max Deg-$d$-Cut is similar. Similar as the proof of previous theorem, we have the following two cases.
    Case (i):
    First, we discuss a trivial case when $G$ has at most $k^{d}$ vertices. Since $G$ has at most $k^{d}$ vertices, the kernelization algorithm returns the same instance $(G,S,k)$. Given a $d$-cut $F^{\prime}$ of $H$, the solution-lifting algorithm just outputs $F=F^{\prime}$. Clearly, $F$ is a $d$-cut of $G$ and this clearly satisfies (i) and (ii) of Definition 2.1. This gives us a polynomial-delay enumeration kernelization with at most $k^{d}$ vertices.
    Case (ii):
    Otherwise, $G$ has more than $k^{d}$ vertices and let $(G,S,k)$ be the input instance. The kernelization algorithm invokes the marking scheme ${\sf MarkVC}(G,S,k)$ as described earlier this section and constructs the graph $H$ such that $S_{0}\cup S$ is a vertex cover of $H$. It follows from Observation 3.1 that $H$ has $\mathcal{O}(d^{3}k^{d+1})$ vertices such that $|S_{0}\cup S|$ is $\mathcal{O}(d^{3}k^{d})$. Output the instance $(H,S_{0}\cup S,|S_{0}\cup S|)$. This completes the description of the kernelization algorithm.
    The solution-lifting algorithm exploits the proof ideas of the lemmas of previous sections, and given a $d$-cut of $H$, constructs equivalent $d$-cuts of $G$ with polynomial-delay as follows.
    Let $F^{*}=E(A^{*},B^{*})$ be a $d$-cut of $H$ and $L({\mathcal{Y}})=\cup_{T\in{\mathcal{Y}}}L_{T}$. Therefore, we infer from Lemma 3.1 that $F^{*}$ is also a $d$-cut of $\widetilde{H}$. Furthermore, every component $C$ of $H[S_{0}]$ is either contained in $A^{*}$ or contained in $B^{*}$. If $F^{*}\cap L({\mathcal{Y}})=\emptyset$, then $F^{*}$ is a $d$-cut of $\widetilde{H}-I({\mathcal{Y}})$. Then, due to Lemma 3.2, $F^{*}$ is a $d$-cut of $G-I({\mathcal{Y}})$. Since $F^{*}$ is equivalent to itself and $F^{*}$ is a $d$-cut of $\widetilde{H}$, it follows from Lemma 3.3 that $F^{*}$ is a $d$-cut of $G$ as well. The solution lifting algorithm just outputs $F^{*}$ as a $d$-cut to the input instance $(G,S,k)$.
    Otherwise it must be that $F_{b}=F^{*}\cap L({\mathcal{Y}})\neq\emptyset$. Let $F_{g}=F^{*}\setminus F_{b}$. We use the following procedure to compute a collection of $d$-cuts of $G$ with $n^{\mathcal{O}(1)}$-delay. The algorithm is recursive and takes parameter as $(J,{\mathcal{D}})$ as follows.
    In order to satisfy the Property (ii) of Definition 2.1, the solution-lifting algorithm keeps $F_{g}$ unchanged. Initialize $J=F_{g}$ and ${\mathcal{D}}={\mathcal{Y}}$. In other words, if $T\in{\mathcal{D}}$, then $T$ is a bad subset of $S$ in $G$ such that $F^{*}\cap L_{T}\neq\emptyset$.
    1. (a)
      If ${\mathcal{D}}=\emptyset$, then output $J$.
    2. (b)
      Else, take an arbitrary $T\in{\mathcal{D}}$ and let $F_{T}=F^{*}\cap L_{T}$. Note that $T$ is a bad subset of $S$ in $G$. Observe that every $uv\in F_{T}$ has one endpoint $u\in T$ and the other endpoint $v\in I_{d}$ such that $N_{\widetilde{H}}(v)=T$. Suppose $P\subseteq I_{d}$ such that $E(P,T)=F_{T}$ and $|P|=\beta$. For every $w\in P$, we look at the number of vertices in a pendant clique $C_{w}$ attached to $w$ such that $C_{w}\setminus\{w\}$ is a connected component of $H[S_{0}]$. Let $b_{w}$ denotes the number of vertices in $C_{w}\setminus\{w\}$. Formally, let $Q_{P}=\{b_{w}\mid w\in P\}$. If $Q_{P}\neq\{2d+1,\ldots,2d+\beta\}$, then we recursively call the algorithm with parameter $J\cup F_{T}$ and ${\mathcal{D}}\setminus\{T\}$. Otherwise, perform the following steps.
    3. (c)
      We are now in the case that $Q_{P}=\{2d+1,\ldots,2d+\beta\}$. In addition, we assume without loss of generality that $T\subseteq A^{*}$ and then $P\subseteq B^{*}$ such that $|P|=\beta$. It means that there are $\beta$ vertices in $B^{*}\cap I_{d}$ whose neighborhood equals exactly $T$. Let $\widehat{I}_{T}=\{u\in V(G)\cap I_{d}\mid N(u)=T\}$, i.e. the set of all vertices of $G-S$ whose neighborhood equals $T$. For every set $X^{\prime}$ of $\beta$ vertices from $\widehat{I}_{T}$ such that $(V(G)\setminus V(H))\cap X^{\prime}\neq\emptyset$ or $X^{\prime}=P$, consider $E(X^{\prime},T)=\{uv\mid u\in X^{\prime},v\in T\}$. For every such set $X^{\prime}$ of $\beta$ vertices from $\widehat{I_{T}}$, recursively call the algorithm with parameter $J\cup E(X^{\prime},T)$ and ${\mathcal{D}}\setminus\{T\}$.
    This completes the description (see Algorithm 1 for more details) of the solution lifting algorithm. Given a $d$-cut $F^{*}=E(A^{*},B^{*})$ of $H$, first, we claim that the solution lifting algorithm (or Algorithm 1) runs in $n^{\mathcal{O}(1)}$-delay and outputs a collection of $d$-cuts of $G$ that are equivalent to $F^{*}$. It follows from Lemma 3.1 that $F^{*}$ is a $d$-cut of $\widetilde{H}$. Note that only the edge set $F_{b}$ is replaced by a collection of edge sets. Let $T\in{\mathcal{Y}}$ be a bad set in $G$ such that $F^{*}\cap L_{T}\neq\emptyset$. Since $T$ is a bad subset of $S$ in $G$ and $F^{*}$ is a $d$-cut of $H$, it follows from Observation 3.3 that $T$ is occupied by $F^{*}$. Assume without loss of generality that $T\subseteq A^{*}$. Then, let $P\subseteq\widehat{I}_{T}\cap B^{*}$ such that $E(P,T)=F^{*}\cap L_{T}$ and $|P|=\beta$. Algorithm 1, that is, either the solution lifting algorithm does not change $P$ at all, or it enumerates a collection of subsets of $\widehat{I}_{T}$ having $\beta$ vertices each. Since $\beta=|P|$, for any new set $F$ in the collection of edge sets, it follows that for any bad subset $T\in{\mathcal{Y}}$ that $|F\cap L_{T}|=|F^{*}\cap L_{T}|$. Moreover, any bad subset was occupied by $F^{*}$ and remains occupied by $F$. Since the set of edges from $F_{g}$ are unchanged, $F$ is equivalent to $F^{*}$. It follows from Lemma 3.3 that $F$ is a $d$-cut of $G$. If ${\mathcal{D}}\neq\emptyset$, the algorithm picks a set $T\in{\mathcal{D}}$ and computes a set of pairwise false twins $P\subseteq I_{d}$ with neighborhood $T$. Observe that $|P|=\beta\leq d$. Notice that there are $\beta$ pendant cliques $C_{w}$ attached to every $w\in P$ such that $C_{w}\setminus\{w\}$ is a connected component of $H[S_{0}]$. If the sizes of these pendant cliques are not exactly $\{2d+2,\ldots,2d+\beta+1\}$, then the algorithm recursively calls with parameter $J\cup E(P,T)$ and ${\mathcal{D}}\setminus\{T\}$. Otherwise, the sizes of these pendant cliques are exactly $\{2d+2,\ldots,2d+\beta+1\}$. Then, the solution lifting algorithm chooses every set $X^{\prime}$ of $\beta$ vertices from $\widehat{I}_{T}$ one by one such that $X^{\prime}\cap V(G-H)\neq\emptyset$ including $X^{\prime}=P$, it constructs a unique set of edges $E(X^{\prime},T)$. Then, it recursively calls the algorithm with parameters $J\cup E(X^{\prime},T)$ and ${\mathcal{D}}\setminus\{T\}$. Note that $|X^{\prime}|=\beta\leq d$ as $T\subseteq A^{*}$. This recursive algorithm enumerates the subsets of $\widehat{I}_{T}$ with $\beta$ vertices, and the depth of the recursion is $|{\mathcal{D}}|$. Hence, this is a standard backtracking enumeration algorithm with $n^{\beta}$ branches, and the depth of the recursion is at most $k^{d}$ as ${\mathcal{D}}\subseteq{{S}\choose{\leq d}}$. The $d$-cuts are outputted only in the leaves of this recursion tree when ${\mathcal{D}}=\emptyset$. Hence, the delay between two consecutive distinct $d$-cuts is at most $k^{d}$. As $k^{d}<n$ and every internal node involves $(k^{d})^{\mathcal{O}(1)}$-time computation, it follows that the delay between outputting two distinct equivalent $d$-cuts is $n^{\mathcal{O}(1)}$.
    It remains to justify that given two distinct $d$-cuts $F^{1}$ and $F^{2}$ of $H$, the collection of $d$-cuts $\mathcal{F}^{1}$ and $\mathcal{F}^{2}$ respectively outputted by the solution-lifting algorithm are pairwise disjoint. Since by Lemma 3.1, $F^{1}$ and $F^{2}$ are $d$-cuts of $\widetilde{H}$, let $(A_{1},B_{1})$ and $(A_{2},B_{2})$ be the bipartitions of the $d$-cuts $F^{1}$ and $F^{2}$ respectively. By definition, $F^{1}=F^{1}_{b}\uplus F^{1}_{g}$ such that $F^{1}_{b}$ contains the intersection of $F^{1}$ with all the bad subsets of $S$, and $F^{1}_{g}$ contains the intersection of $F^{1}$ with all the good subsets of $S$ and $F^{1}\setminus L$. By similar definition, $F^{2}=F^{2}_{b}\uplus F^{2}_{g}$. If $F^{1}_{g}\neq F^{2}_{g}$, then the solution-lifting algorithm does not change the edges from $F^{1}_{g}$ and $F^{2}_{g}$ respectively. Hence, any $F_{1}\in\mathcal{F}^{1}$ will be different from $F_{2}\in\mathcal{F}^{2}$.
    If $F^{1}_{b}\neq F^{2}_{b}$ but $F^{1}_{g}=F^{2}_{g}$, then there is a bad subset $T\in{{\mathcal{Y}}}$ such that $F^{1}\cap L_{T}\neq F^{2}\cap L_{T}$. As $T$ is a bad subset of $S$, due to Observation 3.3, $T$ is occupied by $F^{1}$ as well as by $F^{2}$. We focus on the vertices $u\in I_{d}\cap V(\widetilde{H})$ whose neighborhood is exactly $T$ and $J^{u}_{T}\subseteq F^{1}$. Let $P_{1}$ be those vertices and $E(P_{1},T)=F^{1}\cap L_{T}$. Similarly, we focus on the vertices $u\in I_{d}\cap(V\widetilde{H})$ whose neighborhood is exactly $T$ and $J^{u}_{T}\subseteq F^{2}$. Let $P_{2}$ be those vertices and $E(P_{2},T)=F^{2}\cap L_{T}$. The Algorithm 1 looks at the set of pendant cliques that are attached to the vertices of $P_{1}$ and to the vertices of $P_{2}$. Since $E(P_{1},T)\neq E(P_{2},T)$, then the sizes of the pendant cliques attached to the vertices of $P_{1}$ and to the vertices of $P_{2}$ have different sizes. If the sizes of the cliques attached to the vertices of $P_{1}$ are $\{2d+2,\ldots,2d+|P_{1}|+1\}$, then the solution-lifting algorithm (Algorithm 1) enumerates equivalent edge sets $E(Q_{1},T)$ such that $Q_{1}$ intersects $E(G)\setminus E(H)$. Then the collections of edge sets enumerated by the solution-lifting algorithm are also different. Hence, any $d$-cut of $F^{1}$ is different from any $d$-cut of $F^{2}$. This completes the proof for Enum Deg-$d$-Cut parameterized by the size of vertex cover.
    The kernelization algorithm and the solution-lifting algorithm for Enum Max Deg-$d$-Cut are exactly the same as that of Enum Deg-$d$-Cut. Therefore, by Lemma 3.4, it is ensured that Algorithm 1 outputs equivalent maximal $d$-cuts for Enum Max Deg-$d$-Cut. By arguments similar to Enum Deg-$d$-Cut, we can prove that the Property (ii) is also satisfied and the enumeration algorithm runs in $n^{\mathcal{O}(1)}$-delay.
    Since the above cases are mutually exhaustive, this completes the proof of the second statement.
This completes the proof of the theorem. ∎
1 EnumEqv($J$, ${\mathcal{D}}$)
[…]
3 if ${\mathcal{D}}=\emptyset$ then
4       Output $J$;
5
6for every $T\in{\mathcal{D}}$ do
7       $P\leftarrow$ the vertices of $I_{d}$ whose neighborhood is $T$ and $E(P,T)=F^{*}\cap L_{T}$;
8       Let $|P|=\beta$;
9       $C_{w}\leftarrow$ the pendant clique attached to $w$ such that $C_{w}\setminus\{w\}$ is a component of $H[S_{0}]$;
10       Let $b_{w}\leftarrow|C_{w}\setminus\{w\}|$ and $Q_{P}=\{b_{w}\mid w\in P\}$;
11       if If $Q_{P}\neq\{2d+1,\ldots,2d+\beta\}$ then
12             EnumEqv($J\cup E(P,T)$, ${\mathcal{D}}\setminus\{T\}$);
13
14      else
15             $\widehat{I_{T}}=\{u\in I_{d}\mid N_{G}(u)=T\}$;
16             Recall $P$, the collection of $\beta$ vertices from $\widehat{I_{T}}$;
17             for every $X^{\prime}\in{\widehat{{I_{T}}}\choose{\beta}}$ such that $X^{\prime}=P$ or $X^{\prime}\cap V(G-H)\neq\emptyset$ do
18                   $E(X^{\prime},T)=\{uv\mid u\in X^{\prime},v\in T\}$;
19                   EnumEqv($J\cup E(X^{\prime},T)$, ${\mathcal{D}}\setminus\{T\}$);
20
21
22
Algorithm 1 Solution Lifting Algorithm
4 Enumeration Kernelizations Parameterized by Neighborhood Diversity
In this section, we consider the neighborhood diversity (nd) of the input graph as the parameter. We assume that the neighborhood diversity of the input graph $G$ is $k$, i.e. ${\textsf{nd}}(G)=k$. Let $G$ be a graph, a neighborhood decomposition with minimum number of modules can be obtained in polynomial-time [23]. Let ${\mathcal{U}}=\{X_{1},\ldots,X_{k}\}$ be a minimum-size neighborhood decomposition of $G$. A module $X$ is an independent module if $G[X]$ is an independent set and is a clique module if $G[X]$ is a clique. Let ${\mathcal{G}}$ be the quotient graph for ${\mathcal{U}}$, i.e. $\{X_{1},\ldots,X_{k}\}$ is the set of nodes of ${\mathcal{G}}$. Two distinct nodes $X_{i}$ and $X_{j}$ are adjacent in ${\mathcal{G}}$ if and only if a vertex of the module $X_{i}$ and a vertex of the module $X_{j}$ are adjacent to each other in $G$. Note that if a vertex of $X_{i}$ is adjacent to a vertex in $X_{j}$, then due to the property of module, every vertex of $X_{i}$ is adjacent to every vertex of $X_{j}$. A module is called trivial if it is an independent module and is not adjacent to any other module. Formally, a module $X$ of $G$ is a trivial module if $X$ is an independent set of $G$ and $N_{G}(X)=\emptyset$. A module is called nice if it is an independent module, has at most $d$ neighbors outside and not a trivial module. Formally, $X$ is a nice module of $G$ if $X$ is independent, $|N_{G}(X)|\leq d$ and $X$ is not trivial. Observe that there can be at most one trivial module in $G$. Also, a trivial module is an isolated node in the quotient graph ${\mathcal{G}}$. If a module $X$ is nice, then it is adjacent to at most $r\leq d$ modules $X_{1},\ldots,X_{r}$ such that $\sum_{i=1}^{r}|X_{i}|\leq d$. If a nice module $X$ is adjacent to exactly $r$ modules $X_{1},\ldots,X_{r}$, we say that $\{X_{1},\ldots,X_{r}\}$ is a small module-set. If $r=1$, then for simplicity, we say that $X_{1}$ is a small module. The set of vertices $\cup_{i=1}^{r}X_{i}$ that is spanned by this small module-set $\{X_{1},\ldots,X_{r}\}$ is called a small subset. Let $(G,{\mathcal{U}},k)$ denotes the input instance. We perform the following procedure ${\sf MarkND}(G,{\mathcal{U}},k)$ that is exploited by all the kernelizations in this section.
Procedure ${\sf MarkND}(G,{\mathcal{U}},k)$:
  • Initialize $H^{*}:=G$ and $S_{0}:=\emptyset$.
  • If $X$ is a trivial module of $H^{*}$, then mark an arbitrary vertex of $X$.
  • Let $X$ be a nice module of $G$. If $X$ has at most $23d+$ 2 vertices, then mark all vertices of $X$. Otherwise, if $X$ has more than $23d+$ 2 vertices, then mark $23d+$ 2 vertices of $X$.
  • If a nice module $X$ has $23d+$ 2 marked vertices but $|X|>3d+2$, then choose arbitrary $d$ marked vertices $u_{1},\ldots,u_{d}\in X$. For every $i\in[d]$, attach a pendant clique $C_{i}$ with $12d+i+$ 1 vertices into $u_{i}$ and set $S_{0}:=S_{0}\cup(C_{i}\setminus\{u_{i}\})$. Replace the module $X$ by $X\setminus\{u_{1},\ldots,u_{d}\}$. Additionally, for every $i\in[d]$, create modules $\{u_{i}\}$ and $C_{i}\setminus\{u_{i}\}$.
  • If a module $X$ of $G$ is neither trivial nor nice, then mark $\min\{2d+1,|X|\}$ vertices of $X$.
Let $H$ be the graph constructed by using the vertices of $H$ and theafter invoking the above mentioned procedure ${\sf MarkND}(G,{\mathcal{U}},k)$. Then, the following observation holds true.
Observation 4.1.
Let $G$ be the input graph with a neighborhood decomposition $\{U_{1},\ldots,U_{k}\}$ having minimum number of modules and let $H$ be the graph constructed after invoking ${\sf MarkND}(G,{\mathcal{U}},k)$. Then, the neighborhood diversity of $H$ is at most $(2d+1)k$.
Proof.
Let the premise of the statement be true. Then, ${{\rm nd}}(G)=k$. Let $U_{i}$ be a nice module of $G$ such that $|U_{i}|>3d+2$. By construction, $23d+$ 2 vertices from $U_{i}$ are marked and there are $d$ vertices $u_{1},\ldots,u_{d}\in U_{i}$ such that for every $i\in[d]$, a pendant clique $C_{i}$ of size $12d+i+$ 1 containing $u_{i}$ is attached to it. Observe that the newly constructed modules are $C_{i}\setminus\{u_{i}\}$ and $Y_{i}=\{u_{i}\}$. Hence, this procedure MarkND($G,{\mathcal{U}},k$) constructs at most $2dk$ extra modules. Therefore the total number of modules in $H$ is at most $(2d+1)k$ concluding that ${\textsf{nd}}(H)\leq(2d+1)k$. ∎
[…]
Lemma 4.1.
This lemma has two parts. (i) If $X$ be a module in $G$ with at least $12d+$ 1 vertices that has at least two neighbors outside, then for any $d$-cut $(A,B)$ of $G$, either $N_{G}(X)\subseteq A$ or $N_{G}(X)\subseteq B$. Moreover, if $X$ has at least $d+1$ neighbors outside, then for any $d$-cut $(A,B)$ of $G$, $N_{G}[X]\subseteq A$ or $N_{G}[X]\subseteq B$. (ii) If $X$ is a clique module of $G$ and $|X|>2d$, then for any $d$-cut $(A,B)$ of $G$, either $N_{G}[X]\subseteq A$ or $N_{G}[X]\subseteq B$.
The first part of the above lemma statement is a refinement of Lemma 2.1 and the second part of the above lemma is a generalization of the Observation 1 from [15].
We also prove the following lemma the proof of which is similar to that of Lemma 3.1.
Lemma 4.2.
A set of edges $F\subseteq E(H-S_{0})$ is a $d$-cut of $H-S_{0}$ if and only if $F$ is a $d$-cut of $H$.
Proof.
For the forward direction ($\Rightarrow$), let $F\subseteq E(H-S_{0})$ be a $d$-cut of $H-S_{0}$ and $(A^{\prime},B^{\prime})$ be the vertex bipartition of the set $V(H-S_{0})$ into two parts. Consider every connected component of $H[S_{0}]$. We initialize $A:=A^{\prime}$ and $B:=B^{\prime}$. If $C$ is a connected component of $H[S_{0}]$, then $|C|\geq 2d+1$. If $N_{H}(C)=\{u\}$, then $C\cup\{u\}$ is a pendant clique with articulation point $u$. If $u\in A^{\prime}$, then we set $A:=A\cup C$. Similarly, if $u\in B^{\prime}$, then we set $B:=B\cup C$. This construction of $(A,B)$ maintains the fact that $E(A,B)=E(A^{\prime},B^{\prime})$. Hence, $F$ is a $d$-cut of $H$.
For the backward direction ($\Leftarrow$), let $F=(A,B)$ be a $d$-cut of $H$. Let $C$ be a pendant clique containing at least $22d+$ 2 vertices attached a vertex $u$ such that $C\setminus\{u\}$ is a connected component of $H[S_{0}]$. As $C\setminus\{u\}$ also has at least $12d+$ 1 vertices, it follows from (i) of Lemma 2.1 that $C\subseteq A$ or $C\subseteq B$. Hence, for any edge $uv\in E(H)$ with $u,v\in S_{0}$ or $u\in S_{0},v\in V(H-S_{0})$, it follows that $uv\notin F$. Therefore, $F$ is a $d$-cut of $H-S_{0}$. ∎
Let ${\mathcal{X}}$ denotes the set of all nice modules and trivial modules of $G$ and $V_{{\mathcal{X}}}=\cup_{X\in{\mathcal{X}}}X$. We prove the following lemma using Lemma 4.1.
Lemma 4.3.
Let ${\mathcal{X}}$ denotes the set of all nice modules of $G$ and $V_{{\mathcal{X}}}=\bigcup\limits_{X\in{\mathcal{X}}}X$. A set of edges $F\subseteq E(G-V_{{\mathcal{X}}})$ is a $d$-cut of $G$ if and only if $F$ is a $d$-cut of $H-(S_{0}\cup V_{{\mathcal{X}}})$.
Proof.
The proof arguments are similar to that of the proof of Lemma 3.2. Let $H^{\prime}=H-(S_{0}\cup V_{{\mathcal{X}}})$ and $G^{\prime}=G-V_{{\mathcal{X}}}$.
First, we give the backward direction ($\Leftarrow$) of the proof. Let $(A^{\prime},B^{\prime})$ be a $d$-cut of $H^{\prime}$ and $F=E_{H^{\prime}}(A^{\prime},B^{\prime})$. We prove the existence of a $d$-cut $(A,B)$ of $G^{\prime}$ such that the set of crossing edges of $(A,B)$ and $(A^{\prime},B^{\prime})$ are the same. Initialize $A:=A^{\prime}$ and $B:=B^{\prime}$. Consider a vertex $u\in V(G^{\prime})\setminus V(H^{\prime})$. Then, $u\in X$ such that $X$ is a module in $G^{\prime}$ with at least $12d+$ 1 vertices and $|X\cap V(H^{\prime})|=2d+1$. If $X$ is an independent module, then $N_{G^{\prime}}(X)$ has at least $d+1$ vertices. In particular, $N_{H^{\prime}}(X\cap V(H^{\prime}))$ also has at least $d+1$ vertices. Then, due to Lemma 4.1, it follows that $N_{H^{\prime}}[X\cap V(H^{\prime})]\subseteq A^{\prime}$ or $N_{H^{\prime}}[X\cap V(H^{\prime})]\subseteq B^{\prime}$. If $N_{H^{\prime}}[X\cap V(H^{\prime})]\subseteq A^{\prime}$, then set $A:=A\cup\{u\}$. Similarly, if $N_{H^{\prime}}[X\cap V(H^{\prime})]\subseteq B^{\prime}$, then set $B:=B\cup\{u\}$. We repeat this procedure for every $u\in X$ such that $X$ is a module in $G^{\prime}$ with more than $12d+$ 1 vertices. Observe that this above procedure ensures that the edges crossing the cut $(A,B)$ remains the same as the edges crossing the cut $(A^{\prime},B^{\prime})$. On the other hand, if $X$ is a clique module, then $|X|\geq 2d+1$. Hence, $N_{G^{\prime}}[X]$ as well as $N_{H^{\prime}}[X]$ have more than $2d$ vertices. Again due to part (ii) of Lemma 4.1, $N_{H^{\prime}}[X\cap V(H^{\prime})]\subseteq A^{\prime}$ or $N_{H^{\prime}}[X\cap V(H^{\prime})]\subseteq B^{\prime}$. If $N_{H^{\prime}}[X\cap V(H^{\prime})]\subseteq A^{\prime}$, then set $A:=A\cup\{u\}$. Otherwise, $N_{H^{\prime}}[X\cap V(H^{\prime})]\subseteq B^{\prime}$ and set $B:=B\cup\{u\}$. Similarly, this procedure is repeated for every $u\in X$ such that $X$ is a clique module of $G^{\prime}$ with more than $12d+$ 1 vertices. Similar as before, observe that the set of edges crossing the cut $(A,B)$ and the edges crossing the cut $(A^{\prime},B^{\prime})$ remains the same. Therefore, $F$ is a $d$-cut of $G^{\prime}$.
Let us give the forward direction ($\Rightarrow$) of the proof. Let $(A,B)$ be a $d$-cut of $G^{\prime}$ such that $F=E_{G^{\prime}}(A,B)$. If $F$ is also a $d$-cut of $H^{\prime}$, then we are done. Let $uv\in F$ such that $uv\notin E(H^{\prime})$. Either $u,v\in N_{G^{\prime}}[X]$ for some clique module $X$ in $G^{\prime}$, or $u\in X$ such that $X$ is an independent module of $G^{\prime}$ and $u\in X\cap(V(G^{\prime})\setminus V(H^{\prime}))$. In either case, the module $X$ of $G^{\prime}$ has at least $22d+$ 2 vertices and hence at least $12d+$ 1 vertices in $H^{\prime}$. If $u,v\in N_{G^{\prime}}[X]$ for some clique module $X$ in $G^{\prime}$, then due to part (ii) of Lemma 4.1, it follows that $N_{G^{\prime}}[X]\subseteq A$ or $N_{G^{\prime}}[X]\subseteq B$. Since $u,v\in N_{G^{\prime}}[X]$ implying $u,v\in A$ or $u,v\in B$, it contradicts the assumption that $uv\in F$. On the other hand, if $X$ is an independent module of $G^{\prime}$ and $u\in X\cap(V(G^{\prime})\setminus V(H^{\prime}))$, then $N_{G^{\prime}}(X)$ has at least $d+1$ vertices, because no independent module of $G^{\prime}$ is nice. Then, it follows from Lemma 4.1 that $N_{G^{\prime}}[X]\subseteq A$ or $N_{G^{\prime}}[X]\subseteq B$. Then also $u,v\in N_{G^{\prime}}[X]$ implying $u,v\in A$ or $u,v\in B$ contradicting the assumption that $uv\in F$. Hence, if $F$ is a $d$-cut of $G^{\prime}$, then $F$ is also a $d$-cut of $H^{\prime}$. ∎
Let $X$ be a nice module and $Y_{1},\ldots,Y_{r}$ be the collection of modules that are adjacent to $X$. Then, $\{Y_{1},\ldots,Y_{r}\}$ is a small module-set and $T_{X}=\cup_{i=1}^{r}Y_{i}$ containing at most $d$ vertices is a small subset. We call a small subset $T_{X}$ bad if $|X|\geq 3d+3$. Otherwise, we say that $T_{X}$ is good. Informally speaking, a small subset $T$ is good if and only if $|T|\leq d$ and there exists a nice module $X$ with at most $23d+$ 2 vertices in $G$ such that $N_{G}(X)=T$. Similarly, a small subset $T$ is bad if and only if $|T|\leq d$ and there exists a nice module $X$ with more than $23d+$ 2 vertices in $G$ such that $N_{G}(X)=T$. Let $T$ be a small subset and $X$ be a nice module such that $N_{G}(X)=T$. For every $u\in X$, we focus on the set $J_{T}^{u}=\{uv\in E(G)\mid v\in T\}$. Informally, for a vertex $u$ in a nice module $X$ and a small subset $T$, the edge set $J^{u}_{T}$ denotes the set of all edges one endpoint of which is $u$ and other endpoint is in $T$. Let $T$ be a small subset that is bad. Then, $T$ is said to be broken by an edge set $F$ if there exists $u\in X$ for a nice module $X$ satisfying $N(X)=T$ such that $J^{u}_{T}\cap F\neq\emptyset$ and $J_{T}^{u}\not\subset F$. If a small subset $T$ that is bad is not broken by $F$, then $T$ is said to be occupied by $F$. Equivalently, observe that if $T$ is occupied by $F$, then for every $u\in X$ such that $N(X)=T$, it must be that either $J^{u}_{T}\cap F=\emptyset$ or $J^{u}_{T}\subseteq F$. The following observation holds true the proof of which is similar to that of Observation 3.3.
Observation 4.2.
Let $F\subseteq E(G)$ be a $d$-cut of $G$. Then, if a small subset $T$ is bad, then $T$ is occupied by $F$.
Given a small subset $T$, we use $E_{R}(T)$ to denote the edges of $G$ that has one endpoint in $T$ and the other endpoint in a nice module $Y$ of $G$ such that $N_{G}(Y)=T$. Let ${\mathcal{D}}$ denotes the collection of all small subsets that are bad. Two distinct subsets of edges $F_{1}$ and $F_{2}$ are equivalent if
  • for every $T\in{{\mathcal{D}}}$, $|F_{1}\cap E_{R}(T)|=|F_{2}\cap E_{R}(T)|$,
  • $F_{1}\setminus(\bigcup\limits_{T\in{{\mathcal{D}}}}E_{R}(T))=F_{2}\setminus(\bigcup\limits_{T\in{\mathcal{D}}}E_{R}(T))$, and
  • for every $T\in{{\mathcal{D}}}$, $T$ is occupied by $F_{1}$ if and only if $T$ is occupied by $F_{2}$.
Informally, we focus on the edges that have one endpoint in a small subset $T$ that is bad and the other endpoint in a nice module $X$ whose neighborhood exactly equals $T$. In other words, we focus on the set $E(X,T)$. The first condition stated above says that $|F_{1}\cap E(X,T)|=|F_{2}\cap E(X,T)|$. The second condition above says that if we only focus on the edges that are not incident to the nice modules of size more than $23d+$ 2, then their intersection in both $F_{1}$ and $F_{2}$ are precisely the same. The following set of lemmas illustrate the equivalence between the $d$-cuts of $G$ and the $d$-cuts of $H-S_{0}$ since $H-S_{0}$ is a subgraph of $G$. The proofs of the following lemmas use Lemma 4.1 and Observation 4.2 and have arguments that are similar to the proofs of Lemma 3.3 and Lemma 3.4.
Lemma 4.4.
Let $F^{\prime}\subseteq E(H-S_{0})$ be a $d$-cut of $H-S_{0}$ and let $F\subseteq E(G)$. If $F$ is equivalent to $F^{\prime}$, then $F$ is a $d$-cut of $G$.
Proof.
Let $F^{\prime}\subseteq E(H-S_{0})$ be a $d$-cut of $H-S_{0}$ and $F\subseteq E(G)$ such that $F$ is equivalent to $F$. Suppose that $(A^{\prime},B^{\prime})$ be the corresponding vertex bipartition of the $d$-cut such that $F^{\prime}=E_{H-S_{0}}(A^{\prime},B^{\prime})$. Consider the collection ${\mathcal{D}}$ of all small subsets that are bad. For every $T\in{{\mathcal{D}}}$, there is a nice module $X$ containing at least $33d+$ 3 vertices such that $N_{H-S_{0}}(X)=T$. It suffices to prove that every $u\in V(G)$ is incident to at most $d$ edges in $F$. There are four cases that arise. First case is that $u\in X$ for some nice module $X$ such that $|X|\geq 3d+3$. The second case occurs when $u\in X$ for some nice module $X$ but $|X|\leq 3d+2$. The third case occurs when $u\in T$ for some small subset $T$ that is adjacent to a nice module $X$. Finally, the fourth case occurs when none of the above three cases appear. Since $F$ and $F^{\prime}$ are equivalent it must be that $F\setminus(\bigcup\limits_{T\in{\mathcal{D}}}E_{R}(T))=F^{\prime}\setminus(\bigcup\limits_{T\in{\mathcal{D}}}E_{R}(T))$. Let $\tilde{F}=F\setminus(\bigcup\limits_{T\in{\mathcal{D}}}E_{R}(T))$. Consider any $u\in V(G)$.
Case (i):
When $u\in X$ for some nice module $X$ such that $|X|\geq 3d+3$. Note that $N_{G}(u)=N_{G}(X)=T$ for some small subset $T$ that is bad. Given $u\in X$, let $J^{u}_{T}=\{uv\in E(G)\mid v\in T\}$. Precisely $|T|\leq d$ and $J^{u}_{T}$ be the exact set of edges in $G$ that are incident to $u$. Similarly, for any $u\in X\cap V(H-S_{0})$, $J^{u}_{T}$ denotes the exact set of edges in $H-S_{0}$ that are incident to $u$. As $F^{\prime}$ is a $d$-cut of $H-S_{0}$, it follows from Observation 4.2 that $T$ must be occupied by $F^{\prime}$. Hence, for every $u\in X\cap V(H-S_{0})$, $J_{T}^{u}\subseteq F^{\prime}$ or $J_{T}^{u}\cap F^{\prime}=\emptyset$. As $F$ is equivalent to $F^{\prime}$, it must be that $T$ is occupied by $F$ as well. Therefore, for every $u\in X$, either $J^{u}_{T}\subseteq F$ or $J_{u}^{T}\cap F=\emptyset$. In either case, there are at most $d$ edges in $F^{\prime}$ that are incident to $u$. The same property holds that there are at most $d$ edges in $F$ that are incident to $u$.
Case (ii):
When $u\in X$ for some nice module $X$ of $G$ such that $|X|\leq 3d+2$. Observe that $X\subseteq V(H-S_{0})$. Then, any edge $uv\in F^{\prime}$ incident to $u$ contributes to the set $F^{\prime}\setminus(\bigcup\limits_{T\in{\mathcal{D}}}E_{R}(T))$. Since $F$ is equivalent to $F^{\prime}$, it must be that $F\setminus(\bigcup\limits_{T\in{\mathcal{D}}}E_{R}(T))=F^{\prime}\setminus(\bigcup\limits_{T\in{\mathcal{D}}}E_{R}(T))$. Hence, the same $uv\in E(G)$ incident to $u$ contributes to $F\setminus(\bigcup\limits_{T\in{\mathcal{D}}}E_{R}(T))$. Therefore, if $\alpha$ edges are incident to $u$ that are contributing to $F^{\prime}\setminus(\bigcup\limits_{T\in{\mathcal{D}}}E_{R}(T))$, then $\alpha$ edges are incident to $u$ that are contributing to $F\setminus(\bigcup\limits_{T\in{\mathcal{D}}}E_{R}(T))$. Since $\alpha\leq d$, the claim follows.
Case (iii):
Let $u\in X$ such that $X$ is independent but not a nice module of $G$. Then, $N_{G}(u)$ has more than $d$ vertices. If $|X|\geq 2d+1$, then $X\cap V(H-S_{0})$ has at least $12d+$ 1 vertices. Since $X$ has at least $d+1$ neighbors outside and for every other module $Y$, there are $\min\{|Y|,2d+1\}$ marked vertices, $N_{H-S_{0}}(X)$ also has at least $d+1$ vertices. Therefore, by Lemma 4.1, it follows that $N_{H-S_{0}}[X\cap V(H-S_{0})]\subseteq A^{\prime}$ or $N_{H-S_{0}}[X\cap V(H-S_{0})]\subseteq B^{\prime}$. Then, no edge incident to any vertex of $X\cap V(H-S_{0})$ is in $F^{\prime}$. Since $u\in X$ and no edge incident to any vertex of $X\cap V(H-S_{0})$ is in $F^{\prime}$ and $F$ is equivalent to $F^{\prime}$, no edge incident to $u$ also can be in $F$. This completes the justification that there are at most $d$ edges of $F$ that are incident to $u$.
Case (iv):
Let $u\in X$ for some clique module $X$ in $G$. There are two subcases $|X|\geq d+1$ and $|X|\leq d$ and there is a module $Y$ such that $N_{G}(Y)=X$. If $|X|\geq d+1$, then $X$ cannot be part of a small module set. So, the edges incident to the module $X$ contribute to the edge set $F^{\prime}\setminus(\cup_{T\in{{\mathcal{D}}}}E_{R}(T))$ such that ${{\mathcal{D}}}$ is the collection of all small subsets that are bad. Since it follows due to equivalence between $F$ and $F^{\prime}$ that $F^{\prime}\setminus L=F\setminus L$, Moreover, $F^{\prime}$ is a $d$-cut implying that at most $d$ edges of $F^{\prime}$ are incident to $u\in X$. Hence, the same set of at most $d$ edges are incident to $F$ as well.
The other case is $|X|\leq d$. We have two subcases here. The first subcase occurs when $X\subseteq T$ for a small subset $T$ with at most $d$ vertices. Let $Y$ be a module of $G$ such that $N_{G}(Y)=T$. Define $F^{\prime}_{T,u}(Y)$ the edges of $F^{\prime}$ with one endpoint $u$ and the other endpoint in a module $Y$ such that $N_{H-S_{0}}(Y)=T$. Similarly, define $F_{T,u}(Y)$ the edges of $F$ with one endpoint $u$ and the other endpoint in the same module $Y$ such that $N_{H-S_{0}}(Y)=T$. If $Y$ is a clique module, then $F^{\prime}_{T,u}(Y)$ contributes to the edges $F^{\prime}\setminus(\cup_{T\in{{\mathcal{D}}}}E_{R}(T))$ and $F_{T,u}(Y)$ contributes to the edges in $F\setminus(\cup_{T\in{{\mathcal{D}}}}E_{R}(T))$. Since $F$ is equivalent to $F^{\prime}$, it must be that $F^{\prime}\setminus(\cup_{T\in{{\mathcal{D}}}}E_{R}(T))=F\setminus(\cup_{T\in{{\mathcal{D}}}}E_{R}(T))$. It implies that $F^{\prime}_{T,u}(Y)=F_{T,u}(Y)$. Similarly, let $Y$ be an independent module. If $|Y|\leq 3d+2$, then $T$ is a good subset. Similar as before, we can argue that $F^{\prime}_{T,u}(Y)=F_{T,u}(Y)$. Finally, if $|Y|\geq 3d+3$, then $T$ is a small subset that is bad. Since $F^{\prime}$ is equivalent to $F$, it follows that $|F\cap E_{R}(T)|=|F^{\prime}\cap E_{R}(T)|$. Note that $F^{\prime}_{T,u}(Y)$ contributes to the edges $F^{\prime}\cap E_{R}(T)$ and $F_{T,u}(Y)$ contributes to the edges $F\cap E_{R}(T)$. Moreover, it follows from Observation 4.2 that $T$ is occupied by $F^{\prime}$. Since $F^{\prime}$ is equivalent to $F$, $T$ is occupied by $F$ as well. Clearly, if $|F^{\prime}_{T,u}(Y)|=\alpha$, i.e. there are $\alpha$ edges in $F^{\prime}$ that are incident to $u$ with other endpoint in $Y$, then $\alpha$ edges in $F$ that are incident to $u$ with the other endpoint in $Y$. Therefore, $|F^{\prime}_{T,u}(Y)|=|F_{T,u}(Y)|$. By arguments similar to that of Lemma 3.3, it follows that the number of edges in $F^{\prime}$ that are incident to $u$ is the same as the number of edges in $F$ that are incident to $u$. Hence, there are at most $d$ edges in $F$ that are incient to $u$.
Since the above cases are mutually exhaustive, this completes the proof of our lemma. ∎
Lemma 4.5.
Let $F\subseteq E(G)$. Then, $F$ is a $d$-cut (minimal or maximal $d$-cut, respectively) of $G$ if and only if $H-S_{0}$ has a $d$-cut (minimal or maximal $d$-cut, respectively) $F^{\prime}$ that is equivalent to $F$.
Proof.
For the backward direction ($\Leftarrow$), let $F^{\prime}$ be a $d$-cut of $H-S_{0}$. Since $F^{\prime}\subseteq E(G)$ and $F^{\prime}$ is equivalent to itself, it follows from Lemma 4.4 that $F^{\prime}$ is also a $d$-cut of $G$.
The forward direction ($\Rightarrow$) of the proof works as follows. Let $(A,B)$ be a $d$-cut of $G$ such that $F=E_{G}(A,B)$. We construct a $d$-cut $F^{\prime}=E_{H-S_{0}}(A^{\prime},B^{\prime})$ from $F$ such that $F^{\prime}$ is equivalent to $F$.
We initialize $A^{\prime}:=B^{\prime}:=\emptyset$. Suppose that ${\mathcal{U}}=\{X_{1},\ldots,X_{k}\}$ be the neighborhood decomposition of $G$. If $X_{i}$ is a clique and has more than $2d$ vertices, then by (i) of Lemma 2.1, $X_{i}\subseteq A$ or $X_{i}\subseteq B$. If $X_{i}\subseteq A$, then we set $A^{\prime}:=A^{\prime}\cup(X_{i}\cap V(H-S_{0}))$. Otherwise, $X_{i}\subseteq B$ and we set $B^{\prime}:=B^{\prime}\cup(X_{i}\cap V(H-S_{0}))$. If $X_{i}$ is a clique and $|X_{i}|\leq 2d$, then $X_{i}\subseteq V(H-S_{0})$. Then, we set $A^{\prime}:=A^{\prime}\cup(X_{i}\cap A)$ and $B^{\prime}:=B^{\prime}\cup(X_{i}\cap B)$.
If $X_{i}$ is an independent module, then the following cases arise. When $|X_{i}|\geq 2d+1$, at most $33d+$ 3 vertices and at least $12d+$ 1 vertices from $X_{i}$ are marked in $H-S_{0}$. In case $X_{i}$ is not nice (i.e. not a nice module) $N_{G}(X_{i})$ has at least $d+1$ vertices. Then consider $X_{i}$ and $T=N_{G}(X_{i})$. Due to Lemma 4.1, it follows that $N_{G}[X_{i}]\subseteq A$ or $N_{G}[X_{i}]\subseteq B$. Since $X_{i}$ is not nice and is an independent module with at least $12d+$ 1 vertices $|X_{i}\cap V(H-S_{0})|=2d+1$. Therefore, if $N_{G}[X_{i}]\subseteq A$, then we set $A^{\prime}:=A^{\prime}\cup(N_{G}[X_{i}]\cap V(H-S_{0}))$. Similarly, if $N_{G}[X_{i}]\subseteq B$, we set $B^{\prime}:=B^{\prime}\cup N_{G}[X_{i}]\cap V(H-S_{0}))$.
On the other hand, if $|X_{i}|\leq 2d$ and but $X_{i}$ is independent and not a nice module, then $N_{G}(X_{i})$ has at least $d+1$ vertices. Moreover, it is ensured that $X_{i}\subseteq V(H-S_{0})$. We set $A^{\prime}:=A^{\prime}\cup(X_{i}\cap A)$ and $B^{\prime}:=B^{\prime}\cup(X_{i}\cap B)$.
Finally, our last case remains $X_{i}$ is nice module implying that $N_{G}(X_{i})$ has at most $d$ vertices. Let $T=N_{G}(X_{i})$ be the small subset. Observe that $T\subseteq V(H-S_{0})$ since all vertices of $T$ are marked. If $|X_{i}|\leq 3d+3$, then $X_{i}\subseteq V(H-S_{0})$. Then, we set $A^{\prime}:=A^{\prime}\cup(A\cap X_{i})$ and $B^{\prime}:=B^{\prime}\cup(B\cap X_{i})$. If $|X_{i}|\geq 2d+1$, then due to Lemma 4.1, $T\subseteq A$ or $T\subseteq B$. If $T\subseteq A$, set $A^{\prime}:=A^{\prime}\cup T$ and if $T\subseteq B$, then set $B^{\prime}:=B^{\prime}\cup T$. On the other hand, if $|X_{i}|\leq 2d$, then set $A^{\prime}:=A^{\prime}\cup(T\cap A)$ and $B^{\prime}:=B^{\prime}\cup(T\cap B)$. The last case remains when $|X_{i}|>3d+3$. It holds clearly that $T$ is bad. Since $F$ is a $d$-cut of $G$, it follows from Observation 4.2 that $T$ is occupied by $F=E_{G}(A,B)$ implying that $T\subseteq A$ or $T\subseteq B$. If $T\subseteq A$, then set $A^{\prime}:=A^{\prime}\cup T$ and if $T\subseteq B$, then set $B^{\prime}:=B^{\prime}\cup T$. Similarly, if a vertex $u\in X_{i}\cap V(H-S_{0})$ and $u\in A$, then set $A^{\prime}:=A^{\prime}\cup\{u\}$ when $u\in A$. Similarly, if $u\in B$ and $u\in V(H-S_{0})\cap X_{i}$, then set $B^{\prime}:=B^{\prime}\cup\{u\}$. Since $X_{i}\cap V(H-S_{0})$ has at most $33d+$ 3 vertices, we consider a vertex $u\in X_{i}\setminus V(H-S_{0})$. If $T\subseteq A$, then consider the vertices $X_{i}^{B}=X_{i}\cap B$. We replace by $|X_{i}^{B}|$ many distinct vertices from $(X_{i}\cap V(H-S_{0}))\setminus B$ and put them in $B^{\prime}$.
This completes the construction of $A^{\prime}$ and $B^{\prime}$ and hence finally $F^{\prime}=E_{H-S_{0}}(A^{\prime},B^{\prime})$. Every above mentioned procedure ensures that the constructed cut $F^{\prime}$ is equivalent to $F$. Finally, by arguments similar to Lemma 3.4, we can prove that every $u\in A^{\prime}$ (and for every $u\in B^{\prime}$), there are at most $d$ vertices that are in $B^{\prime}$ and neighbors of $A^{\prime}$. Hence, $F^{\prime}$ is a $d$-cut of $H-S_{0}$.
For minimal and maximal $d$-cuts of $G$, the arguments are exactly the same. Similar to the proof of Lemma 3.4, the following is not hard to see. When there are two distinct $d$-cuts $F\subset\hat{F}$ of $G$, the graph $H-S_{0}$ has two distinct $d$-cuts $F^{\prime}$ equivalent to $F$ and $\hat{F}^{\prime}$ equivalent to $F^{\prime}$ satisfying the same proper containment property $F^{\prime}\subset\hat{F}^{\prime}$. ∎
Finally, we need to illustrate a property of the minimal $d$-cuts that can be proved by using Lemma 4.1.
Lemma 4.6.
Let $F\subseteq E(G)$ be a minimal $d$-cut of $G$ and $X$ be a nice module with more than $2d$ vertices such that $N_{G}(X)=T$. If $F\cap E_{G}(T,X)\neq\emptyset$, then $|T|\leq d$ and there is unique $u\in X$ such that $J^{u}_{T}=F$.
Proof.
Suppose that the premise of the statement holds true and $(A,B)$ denote the vertex bipartition for the $d$-cut $F$. By definition, $E_{G}(T,X)$ denotes the set of edges that have one endpoint in $T$ and the other endpoint in $X$. Suppose for the sake of contradiction that $|T|\geq d+1$. Since $|X|\geq 2d+1$ and $T=N_{G}(X)$, it follows from Lemma 4.1 that $N_{G}[X]\subseteq A$ or $N_{G}[X]\subseteq B$. This contradicts the premise that $E_{G}(T,X)\cap F\neq\emptyset$. Hence, $|T|\leq d$. Since $T$ is a small subset and $N_{G}(X)=T$ such that $(A,B)$ is a $d$-cut of $G$, it follows from the Lemma 4.1 that $T\subseteq A$ or $T\subseteq B$. Our premise also includes that $F\cap E(T,X)\neq\emptyset$. Then, there exists $u\in X$ such that $J^{u}_{T}=\{uv\in E(G)\mid v\in T\}\subseteq F$. If there are two distinct vertices $u,w\in X$ such that $J^{u}_{T}\cup J^{w}_{T}\subseteq F$, then observe that $F\setminus J^{w}_{T}$ also defines a $d$-cut of $G$. Since $F\setminus J^{w}_{T}\subsetneq F$, it contradicts the minimality of $F$. Hence, there is unique $u\in X$ such that $J^{u}_{T}=F$ completing the proof of the lemma. ∎
Finally, the main results of this section can be proved using the previous lemmas and observations of this section.
1 EnumEqv($J$, ${\mathcal{D}}$)
[…]
3 if ${\mathcal{D}}=\emptyset$ then
4       Output $J$;
5
6for every $T\in{\mathcal{D}}$ do
7       Let $T=N_{G}(X)$;
8       $P\leftarrow$ the vertices of $X$ whose neighborhood is $T$ and $E(P,T)=F^{*}\cap L_{T}$;
9       Let $|P|=\beta$ and $Q_{P}=\{b_{w}\mid$ the component of $H[S_{0}]$ attached to $w\in P$ has $b_{w}$ vertices}}\}};
10       if If $Q_{P}\neq\{2d+1,\ldots,2d+\beta\}$ then
11             EnumEqv($J\cup E(P,T)$, ${\mathcal{D}}\setminus\{T\}$);
12
13      else
14             Assuming $T\subseteq A^{*}$, $\beta$ is the number of vertices from ${X}\cap B^{*}$;
15             for every $X^{\prime}\in{{X}\choose{\beta}}$ such that $X^{\prime}=P$ or $X^{\prime}\cap V(G-H)\neq\emptyset$ do
16                   $E(X^{\prime},T)=\{uv\mid u\in X^{\prime},v\in T\}$;
17                   EnumEqv($J\cup E(X^{\prime},T)$, ${\mathcal{D}}\setminus\{T\}$);
18
19
20
Algorithm 2 Solution Lifting Algorithm
See 1.2
Proof.
We first illustrarte a few common characteristics for both the proofs. Note that every module is either nice or independendent but not nice or a clique module. So, there are at most ${{\rm nd}}$ many sets are small subsets of $G$. Let ${\mathcal{X}}$ denotes the collection of all nice modules of $G$. For each of the three problems, enumeration kernelization has two parts. One is kernelization algorithm and the other is solution-lifting algorithm. Let $(G,{\mathcal{U}},k)$ denotes the input instance such that ${\mathcal{U}}$ be a neighborhood decomposition of $G$ with the minimum number of modules and let $|{\mathcal{U}}|=k$. The kernelization algorithm invokes the procedure MarkND($G,{\mathcal{U}},k$) and computes $H$. It follows from Observation 4.1 that ${\textsf{nd}}(H)\leq(2d+1)k=(2d+1){\textsf{nd}}(G)$. Moreover, every module of $H$ has at most $3d$ vertices. Hence, the number of vertices in $H$ is $\mathcal{O}(d^{2}k)$.
The solution-lifting algorithms are different for different versions of the problem. In the first item given below, we describe the solution-lifting algorithm for Enum Min Deg-$d$-Cut. Finally, in the second item given below, we desceribe the solution-lifting algorithm for Enum Deg-$d$-Cut and Enum Max Deg-$d$-Cut.
  1. (i)
    The solition-lifting algorithm for Enum Min Deg-$d$-Cut works as follows.
    Let $F^{\prime}$ be a minimal $d$-cut of $H$. It follows from to the proper containment property of Lemma 4.5 that the same edge set $F^{\prime}$ is also a minimal $d$-cut of $H-S_{0}$. Let $(A^{\prime},B^{\prime})$ be the vertex bipartition of $H-S_{0}$ corresponding to the $d$-cut $F^{\prime}$. Let $X\in{\mathcal{U}}$ denotes a module of $G$ and $X^{\prime}=X\cap V(H-S_{0})$ denotes the intersection of $X$ with the vertices of $H-S_{0}$. Clearly, if $X$ is a nice module of $G$ and $|X|\leq 3d+2$, then $X=X^{\prime}$. Similarly, if $X$ is a nice module of $G$ and $|X|>3d+2$, then $|X^{\prime}|=3d+2$. Finally, if $X$ is not a nice module of $G$ and $|X|\leq 2d+1$, then $X=X^{\prime}$ and otherwise ($X$ is not a nice module and $|X|>2d+1$) it follows that $|X^{\prime}|=2d+1$. For every nice module $X\in{\mathcal{U}}$, let $T_{X}=N_{G}(X)$. Consider the set ${\mathcal{T}}=\{N_{G}(X):X$ is a nice module of $G$ and $|X|\geq 3d+3\}$. Informally, if a small subset $T\in{\mathcal{T}}$, then $T=N_{G}(X)$ for a nice module $X$ containing more than $23d+$ 2 vertices. Observe that $T_{X}$ is unchanged in $H$ since $|T_{X}|\leq d$. Consider the edges $E(T_{X},X^{\prime})$ such that $T_{X}\in{\mathcal{T}}$. If $F^{\prime}$ is a minimal $d$-cut and $E(T_{X},X^{\prime})\cap F^{\prime}\neq\emptyset$, it follows from Lemma 4.6 that there is unique $u\in X^{\prime}$ such that $A^{\prime}=\{u\}$ and $B^{\prime}=V(H-(S_{0}\cup\{u\})$. But there are $d$ vertices in $X^{\prime}$ each of which has pendant clique attached in $H$. If $C^{*}$ is a pendant clique containing exactly $22d+$ 2 vertices in $H$ that is attached to $u$ (such that $A^{\prime}=\{u\}$) and $C^{*}\setminus\{u\}$ is a connected component of $H[S_{0}]$, then the solution-lifting algorithm outputs the $d$-cuts $(\hat{A},\hat{B})$ of $G$ such that $\hat{A}=\{v\}$ for all $v\in(X\setminus X^{\prime})\cup\{u\}$ and $\hat{B}=B^{\prime}$. Otherwise either the pendant clique $C^{*}$ attached to $u$ has more than $22d+$ 2 vertices in $H$ such that $C^{*}\setminus\{u\}\subseteq S_{0}$ or no pendant clique is attached to $u$ at all. In such a case, the solution-lifting algorithm just outputs $\hat{A}=\{u\}$ and $\hat{B}=V(G)\setminus\{u\}$ as the only $d$-cut of $G$. Note that in this case, $F=E(\hat{A},\hat{B})=E(A^{\prime},B^{\prime})=F^{\prime}$.
    Similarly, if $E(T_{X},X)\cap F^{\prime}=\emptyset$ for every nice module $X$ of $G$ such that $|X|\geq 3d+3$, then $F^{\prime}\cap(\cup_{T\in{\mathcal{D}}}E_{R}(T))=\emptyset$. Note that $F^{\prime}$ is equivalent to itself. Therefore, by Lemma 4.4, $F^{\prime}$ is also a $d$-cut of $G$. Our solution-lifting algorithm just outputs $F=F^{\prime}$ as the only $d$-cut of $G$. As argued already, the correctness of this part follows from Lemma 4.3. Observe that the above set of situations are mutually exhaustive and it satisfies the property (ii) of fully-polynomial enumeration kernel. Finally observe that this solution-lifting algorithm runs in polynomial-time. Therefore, the property (ii) is also satisfied. Hence, Enum Min Deg-$d$-Cut parameterized by nd admits a fully-polynomial enumeration kernelization with $\mathcal{O}(d^{2}{\textsf{nd}})$ vertices.
  2. (ii)
    We describe the solution-lifting algorithms for Enum Max Deg-$d$-Cut and Enum Deg-$d$-Cut. Same as Theorem 1.1, the solution-ligting algorithm for both the problems are the same.
    Let $F^{\prime}$ be a $d$-cut of $H$. Note that due to Lemma 4.2, it follows that $F^{\prime}$ is a $d$-cut of $H-S_{0}$ as well. Our solution-lifting algorithm works in a similar process to the one in the proof of Theorem 1.1. For every nice module $X$ of $G$, let $T_{X}=N_{G}(X)$. If $|X|>3d+2$, then we put $N_{G}(X)$ into ${\mathcal{D}}$. Observe that every $T\in{\mathcal{D}}$ is a small subset that is bad. Observe that there are at most $k$ distinct small subset of size at most $d$ that are bad. Let $F_{2}=\bigcup\limits_{T_{X}\in{\mathcal{D}}}(F^{\prime}\cap E(T_{X},X))$ and $F^{\prime}=J\cup F_{2}$. Since $F^{\prime}$ is a $d$-cut of $H-S_{0}$, it must be that every $T\in{{\mathcal{D}}}$ is occupied by $F^{\prime}$. Also, for any edge set $F\subseteq E(G)$ that are equivalent to $F^{\prime}$, it holds that for every $T_{X}\in{{\mathcal{D}}}$, $|F^{\prime}\cap E(T_{X},X)|=|F\cap E(T_{X},X)|$. We enumerate the collection of equivalent edge sets as follows. Initialize $J:=F^{\prime}\setminus F_{2}$.
    • If ${\mathcal{D}}=\emptyset$, then observe that $J=F^{\prime}$ and the solution-lifting algorithm outputs $J$ as the only $d$-cut of $G$. It follows from Lemma 4.3 that $J$ is also a $d$-cut of $G$.
    • Otherwise, take an arbitrary $T\in{{\mathcal{D}}}$ and let $F_{T}=E(T_{X},T)\cap F^{\prime}$. Every $uv\in F_{T}$ has one endpoint in $u\in T_{X}$ and other endpoint in $v\in X$ and more over $N_{H-S_{0}}(v)=T_{X}$. Suppose $P\subseteq X$ such that $E(T_{X},P)=F_{T}$.
    • We look at the pendant cliques that are attached vertices of $P$. For every $v\in P$, let $C_{v}$ denotes the pendant clique attached to $v$ such that $C_{v}\setminus\{v\}$ is a connected component of $H[S_{0}]$. We see $b_{v}$, i.e. the number of vertices in $C_{v}\setminus\{v\}$. Let $Q_{P}$ denotes the collection of $b_{v}$ values for every $v\in P$ and $|P|=\beta$. If $Q_{P}\neq\{2d+1,\ldots,2d+\beta\}$, then we set $J:=J\cup F_{T}$ and ${\mathcal{D}}:={\mathcal{D}}\setminus\{T_{X}\}$. Then recursively call the algorithm with $(J,{\mathcal{D}})$. Otherwise, perform the following steps.
    • We are now in the case when $Q_{P}=\{2d+1,\ldots,2d+\beta\}$. Same as before, we assume without loss of generality that $T_{X}\subseteq A^{*}$ and $P\subseteq B^{*}$ such that $|P|=\beta$. For every subset $X^{\prime}$ of $\beta$ vertices from $X$ such that $X^{\prime}\cap(V(G)\setminus V(H))\neq\emptyset$ or $X^{\prime}=P$, consider $E(T_{X},X^{\prime})=\{uv\in E(G)\mid u\in T_{X},v\in X^{\prime}\}$. For every such subset $X^{\prime}$ of $\beta$ vertices from $X$, we recursively call the algorithm with $J\cup E(T_{X},X^{\prime})$ and ${\mathcal{D}}\setminus\{T_{X}\}$.
    This completes the description of the solution-lifting algorithm. By construction it follows that any edge set that is outputted is equivalent to $F^{\prime}$. Since $F^{\prime}$ is a $d$-cut of $H-S_{0}$, it follows from Lemma 4.4 that any edge set that is outputted is also a $d$-cut of $G$. Furthermore, note that this algorithm is recursive and it has nd depth with $n^{\beta}$ branches for every $\beta\in[\textsf{nd}]$. Therefore, the time taken to output two distinct $d$-cuts is $n^{\mathcal{O}(1)}$. Hence, the solution-lifting algorithm runs in polynomial-delay. By arguments similar to Theorem 1.1, we can prove that this solution-lifting algorithm satisfies the propoerty (ii).
This completes the proof of this theorem. ∎
5 Enumeration Kernels Parameterized by the Clique Partition Number
In this section, we consider the clique partition number, i.e. pc of the input graph as the parameter. Since the clique partition number of $G$ is the complement of the chromatic number of the complement graph of $G$, for any $k\geq 3$, it is NP-complete to decide whether ${\textsf{pc}}(G)=k$. Therefore, we assume that a clique partition $\{C_{1},\ldots,C_{k}\}$ of $G$ with the minimum number of cliques is given to us with the input graph. We denote $(G,{\mathcal{C}},k)$ the input instance where ${\mathcal{C}}$ is a clique-partition of $G$ into $k$ cliques. Clearly, by our assumption ${\textsf{pc}}(G)=k$. We call a clique $C\in{\mathcal{C}}$ large if $|C|\geq 2d+1$. If a clique $C\in{\mathcal{C}}$ is not a large clique, we call it a small clique.
Reduction Rule 5.1.
Let $C=\{u_{1},\ldots,u_{i}\}$ be a small clique. Then, replace $C$ by $i$ distinct cliques $\{u_{1}\},\ldots,\{u_{i}\}$.
Observe that the above reduction rule does not change the graph. The collection of cliques in ${\mathcal{C}}$ gets updated and $|{\mathcal{C}}|$ increases to at most $2dk$. We first apply Reduction Rule 5.1 exhaustively. We do not apply this above reduction rule anymore after that. When the above reduction rule is not applicable, it holds that any small clique has just one vertex. The next reduction rule is the following.
Reduction Rule 5.2.
If there are two cliques $C_{i},C_{j}\in{\mathcal{C}}$ such that there is $u\in C_{i}$ having more than $d$ neighbors in $C_{j}$ or there is $u\in C_{j}$ having more than $d$ neighbors in $C_{i}$, then make every vertex of $C_{i}$ adjacent to every vertex of $C_{j}$. Replace $C_{i}$ and $C_{j}$ from ${\mathcal{C}}$ by a new clique $C^{\prime}=C_{i}\cup C_{j}$. Reduce $k$ by 1.
We apply Reduction Rule 5.2 only when Reduction Rule 5.1 is not applicable. When the above reduction rule is applicable for a pair of cliques $C_{i},C_{j}\in{\mathcal{C}}$, then either $C_{i}$ or $C_{j}$ (or both) is (or are) a large clique(s). Observe that for any $d$-cut $(A,B)$ and for any (large) clique $C$, it holds that $C\subseteq A$ or $C\subseteq B$. Hence, it is not possible that $C_{i}\subseteq A$ and $C_{j}\subseteq B$. It means that either $C_{i}\cup C_{j}\subseteq A$ or $C_{i}\cup C_{j}\subseteq B$. We apply Reduction Rules 5.1 exhaustively followed by the Reduction Rule 5.2 exhaustively to obtain the instance $(G_{1},{\mathcal{C}}_{1},k_{1})$. After that we do not apply the above mentioned reduction rules anymore. Let $(G_{1},{\mathcal{C}}_{1},k_{1})$ be the obtained instance after exhaustive application of Reduction Rule 5.2 and let ${\mathcal{C}}_{1}$ be the set of obtained cliques. Note that for every two distinct cliques $C_{i},C_{j}\in{\mathcal{C}}_{1}$, any $u\in C_{i}$ or $u\in C_{j}$ can have at most $d$ neighbors in $C_{j}$ or in $C_{i}$, respectively. But it is possible that some $u\in C_{i}$ may have more than $d$ neighbors outside $C_{i}$. If a vertex $u\in C_{i}$ has at most $d$ neighbors outside $C_{i}$, then we say that $u$ is a low vertex. On the other hand if $u\in C_{i}$ has more than $d$ neighbors outside $C_{i}$, then we say that $u$ is a high vertex.
Lemma 5.1.
A set of edges $F\subseteq E(G)$ is a $d$-cut of $G$ if and only if $F$ is a $d$-cut of $G_{1}$.
Proof.
Let $F=(A,B)$ be a $d$-cut of $G$. Observe that Reduction Rule 5.1 does not change the graph. We apply Reduction Rule 5.2 only after an exhaustive application of Reduction Rule 5.1. Consider two cliques $C,D$ such that $E(C,D)$ is not a $d$-cut. Then, it must be that $D$ (or $C$) is a large clique. It implies that $C\cup D\subseteq A$ or $C\cup D\subseteq B$. Hence, no edge $uv\in E(G_{1})\setminus E(G)$ can be present in a $d$-cut of $G$. It means that $F$ is a $d$-cut of $G_{1}$. By similar argument as above, we can prove that any $d$-cut of $G_{1}$ is also a $d$-cut of $G$. ∎
After we obtain the instance $(G_{1},{\mathcal{C}}_{1},k_{1})$, we invoke a marking procedure ${\sf MarkCP}(G_{1},{\mathcal{C}}_{1},k_{1})$ that works as follows.
  1. 1.
    Suppose there is a clique $C\in{\mathcal{C}}_{1}$ and $i$ distinct large cliques $C_{1},C_{2},\ldots,C_{i}\in{\mathcal{C}}_{1}$ such that $C$ has a low vertex $u$ having neighbors only in each of the cliques $C_{1},\ldots,C_{i}$. Formally, we assume that there is $u\in C$ such that $u$ is a low vertex, $N_{G}(u)\setminus C\subseteq\bigcup\limits_{j=1}^{i}C_{j}$, and $\sum\limits_{j=1}^{i}|N_{G}(u)\cap C_{j}|=\ell\leq d$. Then we mark $u$ and the vertices in $\bigcup\limits_{j=1}^{i}(N_{G}(u)\cap C_{j})$. Repeat this marking scheme for every $i\leq d$ cliques and every $\ell\leq d$.
  2. 2.
    If there is a clique $C\in{\mathcal{C}}_{1}$ and $i$ distinct large cliques $C_{1},C_{2},\ldots,C_{i}\in{\mathcal{C}}_{1}$ such that $C$ has a high vertex $u$ satisfying $\sum\limits_{j=1}^{i}|N_{G}(u)\cap C_{j}|>d$ but $\sum\limits_{j=1}^{i-1}|N_{G}(u)\cap C_{j}|\leq d$ then mark $u$ and all the vertices in $\bigcup\limits_{j=1}^{i}(N_{G}(u)\cap C_{j})$. Let $\bigcup\limits_{j=1}^{i}(N_{G}(u)\cap C_{j})$ has $\ell$ vertices. Informally, if for a clique $C\in{\mathcal{C}}_{1}$ and $i$ distinct large cliques $C_{1},\ldots,C_{i}\in{\mathcal{C}}_{1}$, it happens that $u\in C$ is a high vertex but has at most $d$ cumulative neighbors in $C_{1},\ldots,C_{i-1}$ but has more than $d$ cumulative neighbors in $C_{1},\ldots,C_{i}$, then mark $u$ and all its neighbors from $C_{1},\ldots,C_{i}$. Repeat this process for every $i\leq d+1$ and observe that $\ell\leq 2d$.
Informally, the above marking scheme has marked $u\in C$ and all the neighbors of $u$ outside $C$ in other large cliques if there are at most $d$ distinct neighbors. Otherwise, it has marked $u\in C$ and at most $2d$ neighbors in other large cliques. After performing ${\sf MarkCP}(G_{1},{\mathcal{C}}_{1},k_{1})$, we apply the following reduction rule that removes some unnecessary unmarked vertices from large cliques.
Reduction Rule 5.3.
Let $C$ be a large clique in ${\mathcal{C}}_{1}$ and $X\subseteq C$ be the unmarked vertices of $C$. Remove $\min\{|X|,|C|-2d-1\}$ unmarked vertices from $C$.
Let $(G^{\prime},{\mathcal{C}}^{\prime},k^{\prime})$ be the obtained after applying the marking scheme ${\sf MarkCP}(G_{1},{\mathcal{C}}_{1},k_{1})$ followed by an exhaustive application of Reduction Rule 5.3 on $(G_{1},{\mathcal{C}}_{1},k_{1})$. Let ${\mathcal{C}}^{\prime}$ be the set of cliques such that $\bigcup\limits_{C\in{\mathcal{C}}^{\prime}}V(C)=V(G^{\prime})$. The following observation is simple to see.
Observation 5.1.
Let $F=(A^{\prime},B^{\prime})$ be a $d$-cut of $G^{\prime}$ and ${\mathcal{C}}^{\prime}=\{C_{1}^{\prime},\ldots,C_{r}^{\prime}\}$ such that $r\leq(2d+1)k$. Then, there is a partition $I\uplus J=\{1,\ldots,r\}$ such that $A^{\prime}=\bigcup_{i\in I}C_{i}^{\prime}$ and $B^{\prime}=\bigcup_{j\in J}C_{j}^{\prime}$.
We prove the following two lemmas that are among the critical ones to prove our final theorem statement.
Lemma 5.2.
Let $(G,{\mathcal{C}},k)$ be an instance of Enum Deg-$d$-Cut (or Enum Min Deg-$d$-Cut or Enum Max Deg-$d$-Cut) such that $V(G)$ is partitioned into a set ${\mathcal{C}}$ of $k$ cliques. There is an algorithm that runs in $k^{d+1}n^{\mathcal{O}(1)}$-time and constructs $(G^{\prime},{\mathcal{C}}^{\prime},k^{\prime})$ such that $k^{\prime}\leq k$ and $|V(G^{\prime})|$ is $\mathcal{O}(k^{d+2})$.
Proof.
Let us first describe the algorithm. Given $(G,{\mathcal{C}},k)$, at the first phase, the algorithm performs Reduction Rules 5.1 exhaustively followed by Reduction Rule 5.2 exhaustively to obtain $(G_{1},{\mathcal{C}}_{1},k_{1})$ such that $V(G_{1})$ can be partitioned into a set ${\mathcal{C}}_{1}$ of $k_{1}\leq 2dk$ cliques. In the second phase, the algorithm performs the marking scheme ${\sf MarkCP}(G_{1},{\mathcal{C}}_{1},k_{1})$. In the third phase, the algorithm applies Reduction Rule 5.3 on $(G_{1},{\mathcal{C}}_{1},k_{1})$ exhaustively to obtain $(G^{\prime},{\mathcal{C}}^{\prime},k^{\prime})$ such that $V(G^{\prime})$ is partitioned into a set ${\mathcal{C}}^{\prime}$ of $k^{\prime}$ cliques. By construction, $k^{\prime}=k_{1}\leq 2dk$. Furthermore, the subroutine ${\sf MarkCP}(G_{1},{\mathcal{C}}_{1},k_{1})$ takes $k^{d+1}n^{\mathcal{O}(1)}$-many operations. This completes the description of the polynomial-time algorithm to construct the output instance $(G^{\prime},{\mathcal{C}}^{\prime},k^{\prime})$.
Next, we justify that $|V(G^{\prime})|$ is $\mathcal{O}(k^{d+2})$. Observe that by construction, if $C\in{\mathcal{C}}$ is a small clique of $G$, then $C$ is broken into $|C|$ different small cliques in ${\mathcal{C}}^{\prime}$ and the resultant cliques are small cliques of $G^{\prime}$. Otherwise, consider the case when $C\in{\mathcal{C}}$ is a large clique. The algorithm preserves the fact that either $C$ remains a large clique in ${\mathcal{C}}^{\prime}$ or is replaced by another large clique $C^{\prime}$ in ${\mathcal{C}}^{\prime}$. If we justify that every large clique of ${\mathcal{C}}^{\prime}$ has $\mathcal{O}(k^{d+1})$ vertices, then we are done. Consider an arbitrary large clique $\hat{C}\in{\mathcal{C}}^{\prime}$. For every marked vertex $u\in\hat{C}$, we associate a set ${\mathcal{D}}_{u}$ of at most $d+1$ large cliques from ${\mathcal{C}}^{\prime}\setminus\{\hat{C}\}$ such that $u$ has some neighbors in every $C^{\prime}\in{\mathcal{D}}_{u}$ and all those neighbors are marked neighbors. Observe that for every marked vertex $u\in\hat{C}$, the algorithm has marked at least one and at most $2d$ neighbors $v_{1},\ldots,v_{i}$ ($i\leq 2d$) that are outside $C$. Hence, for every marked vertex $u\in\hat{C}$, we can associate a set ${\mathcal{D}}_{u}$ of at most $d+1$ cliques in ${\mathcal{C}}^{\prime}$. On the other hand, let ${\mathcal{D}}$ be a set of at most $d+1$ large cliques $C_{1},\ldots,C_{i}\in{\mathcal{C}}^{\prime}$ with vertices $u_{j}\in C_{j}$ for $j\leq i$. Then, the marking procedure ${\sf MarkCP}$ marks $d$ neighbors of $u_{j}$ in the clique $\hat{C}$. Hence, the number of marked vertices in $\hat{C}$ is at most $\sum\limits_{i=1}^{d+1}d{{k}\choose{i}}$. It implies that every large clique $\hat{C}\in{\mathcal{C}}^{\prime}$ has $\mathcal{O}(k^{d+1})$ vertices. This completes the justification that $|V(G^{\prime})|$ is $\mathcal{O}(k^{d+2})$. ∎
The following lemma illustrates a bijection between the collection of $d$-cuts of $G^{\prime}$ to the collection of $d$-cuts of $G_{1}$.
Lemma 5.3.
Let ${\mathcal{C}}_{1}=\{\hat{C}_{1},\ldots,\hat{C}_{r}\}$ and ${\mathcal{C}}^{\prime}=\{C_{1}^{\prime},\ldots,C_{r}^{\prime}\}$ be the clique-partition of $G_{1}$ and $G^{\prime}$ respectively and $I\uplus J=\{1,\ldots,r\}$ be a partition of $\{1,\ldots,r\}$. Then, $\hat{F}=(\bigcup\limits_{i\in I}\hat{C}_{i},\bigcup\limits_{j\in J}\hat{C}_{j})$ is a $d$-cut of $G_{1}$ if and only if $F^{\prime}=(\bigcup\limits_{i\in I}C_{i}^{\prime},\bigcup\limits_{j\in J}C_{j}^{\prime})$ is a $d$-cut of $G^{\prime}$.
Proof.
The forward direction ($\Rightarrow$) is easy to see because $G^{\prime}$ is a subgraph of $G_{1}$.
Let us give the proof of the backward direction ($\Leftarrow$). Suppose that for a partition $I\uplus J=\{1,\ldots,r\}$, $(\bigcup\limits_{i\in I}C_{i}^{\prime},\bigcup\limits_{j\in J}C_{j}^{\prime})$ is a $d$-cut of $G^{\prime}$. We assume for the sake of contradiction that $\hat{F}=(\bigcup\limits_{i\in I}\hat{C}_{i},\bigcup\limits_{j\in J}\hat{C}_{j})$ is not a $d$-cut of $G_{1}$. Then there exists $u\in\bigcup\limits_{i\in I}\hat{C}_{i}$ such that $u$ has $d+1$ neighbors in $\bigcup\limits_{j\in J}\hat{C}_{j}$.
Let us consider the case when $u$ is a marked vertex, implying that $u\in\bigcup\limits_{i\in I}C_{i}^{\prime}$. Then, there is $v\in N_{G_{1}}(u)$ such that $v$ is an unmarked neighbor and $v\notin V(G^{\prime})$. Then, there exists $t\in\{1,\ldots,r\}$ such that $v\in\hat{C}_{t}\setminus C_{t}^{\prime}$. Recall that at most $2d$ neighbors of $u$ were marked by the marking procedure. Since there is a neighbor $v\in\hat{C}_{t}$ is unmarked, $u$ had overall at most $d$ neighbors in some other large cliques $\hat{C}_{1},\ldots,\hat{C}_{p}$. It means that $\sum\limits_{\ell=1}^{p}(|N_{G_{1}}(u)\cap\hat{C}_{\ell}|)\leq d$ but $\sum\limits_{\ell=1}^{p}(|N_{G_{1}}(u)\cap\hat{C}_{\ell}|)+|N_{G_{1}}(u)\cap\hat{C}_{t}|>d$. In particular, $u$ is a high vertex. But then more than $d$ marked neighbors of $u$ are present in $G^{\prime}$. As at most $d$ marked neighbors of $u$ are already present in $\bigcup\limits_{j\in J}\hat{C}_{j}$, the same set of marked neighbors of $u$ were also present in $\bigcup\limits_{j\in J}C_{j}^{\prime}$. But, it is also true that there is an unmarked neighbor $v\in\hat{C}_{t}$. But then $u$ has other marked neighbors in $C_{t}^{\prime}$ that are also present in $\bigcup\limits_{j\in J}C_{j}^{\prime}$. But then $u$ has a total of more than $d$ marked neighbors present in $\bigcup\limits_{j\in J}C_{j}^{\prime}$. This leads to a contradiction to the hypothesis that $F^{\prime}$ is a $d$-cut of $G^{\prime}$.
The other case is when $u$ is unmarked vertex of $\hat{C}_{p}$ for some $p\leq r$. Let us consider these $d+1$ neighbors of $u$ that are distributed across $j\geq 2$ distinct cliques from $\{\hat{C}_{i}\mid i\in J\}$. A reason $u$ is unmarked because there is some other marked vertex $\hat{u}\in\hat{C}_{p}$ such that $\hat{u}$ also has at least $d+1$ neighbors in the same set of $j\geq 2$ distinct cliques from $\{\hat{C}_{i}\mid i\in J\}$. As $\hat{u}$ is marked in $G_{1}$, it must be that $\hat{u}\in G^{\prime}$. In fact in $\hat{u}\in\bigcup\limits_{i\in I}C_{i}^{\prime}$ has at least $d+1$ neighbors in $\bigcup\limits_{j\in J}C_{j}^{\prime}$. Then it leads to a contradiction to the hypothesis that $F^{\prime}$ is also not a $d$-cut of $G^{\prime}$. This completes the proof. ∎
The above lemma provides us a proper bijection from the set of all $d$-cuts of $G_{1}$ to all the $d$-cuts of $G^{\prime}$. Using this characterization (Lemma 5.3) and Lemma 5.2, we can prove our result of this section. Kernelization algorithms exploit the procedure MarkCP and the reduction rules above.
See 1.3
Proof.
We prove this for Enum Deg-$d$-Cut. The proof arguments for Enum Max Deg-$d$-Cut and Enum Min Deg-$d$-Cut will be similar. Let $(G,{\mathcal{C}},k)$ be an input instance to Enum Deg-$d$-Cut and $\{C_{1},\ldots,C_{k}\}$ be the set of $k$ cliques satisfying $V(G)=C_{1}\uplus\cdots\uplus C_{k}$. If $|V(G)|\leq k^{d+2}$, then we output $(G,{\mathcal{C}},k)$ as the output instance. So, we can assume without loss of generality that $|V(G)|>k^{d+2}$. We invoke Lemma 5.2 to obtain $(G^{\prime},{\mathcal{C}}^{\prime},k^{\prime})$. Since the algorithm by Lemma 5.2 runs in $k^{d+1}n^{\mathcal{O}(1)}$-time, and we invoke this algorithm only when $n>k^{d+2}$, this reduction algorithm runs in polynomial-time. This completes the description of the reduction algorithm and $|V(G^{\prime})|$ is $\mathcal{O}(k^{d+2})$.
The solution lifting algorithm works as follows. Let $F^{\prime}$ be a $d$-cut of $(G^{\prime},{\mathcal{C}}^{\prime},k^{\prime})$ such that ${\mathcal{C}}^{\prime}$ be the set of cliques for the clique partition of $G^{\prime}$. It follows from Observation 5.1 that for any clique $C^{\prime}\in{\mathcal{C}}^{\prime}$, $E(G^{\prime}[C^{\prime}])\cap F^{\prime}=\emptyset$. Let $I\uplus J=\{1,\ldots,k^{\prime}\}$ be a partition such that $F^{\prime}=E(\bigcup\limits_{i\in I}C_{i}^{\prime},\bigcup\limits_{j\in J}C_{j}^{\prime})$. The solution lifting algorithm first computes $\hat{F}=E(\bigcup\limits_{i\in I}\hat{C}_{i},\bigcup\limits_{j\in J}\hat{C}_{j})$. Due to Lemma 5.3, it holds that $E(\bigcup\limits_{i\in I}\hat{C}_{i},\bigcup\limits_{j\in J}\hat{C}_{j})$ is a $d$-cut of $G_{1}$. After that the solution lifting algorithm outputs $F=\hat{F}$ as the $d$-cut of $G$. Again from Lemma 5.1, $F$ is a $d$-cut of $G$. Clearly, both the reduction algorithm and the solution lifting algorithms run in polynomial-time. Hence, it is a bijective enumeration kernel with $\mathcal{O}(k^{d+2})$ vertices. ∎
6 Conclusions and Future Research
We have initiated the study of (polynomial-delay) enumeration kernelization for variants of Degree-$d$-Cut problem. We have provided (polynomial-delay) enumeration kernels when parameterized by the size of vertex cover, the size of neighborhood diversity and clique partition number of the graph. In addition, we have also explained why we are unlikely to obtain a polynomial-delay enumeration kernel when parameterized by treedepth, cliquewidth etc. The negative result of the Proposition 2.2 can also be extended when we consider a combined parameter – treedepth, the number of edges in the cut and maximum degree of the graph. Obtaining our positive results has required proving some properties that are more specific to $d$-cuts of a graph for any arbitrary $d\geq 2$ in addition to the properties specific to the parameter we consider (see Theorem 7 of [15]). We believe that a lot of follow-up research works can be done for Enum Deg-$d$-Cut and the other enumeration variants. In particular, we have realized that there are some challenges to obtain (polynomial-delay) enumeration kernels when we consider structurally smaller parameters, e.g. feedback vertex set size, cluster vertex deletion set size, modular-width etc. An interesting future research direction would be to explore the above mentioned other structural parameters, the status of which are open even for $d=1$, i.e. for the enumeration variant of the Matching Cut problem. Another interesting research direction would be to identify a parameterized problem the decision version of which admits polynomial kernel but the enumeration version of the same does not. Obtaining such a result (if any) would need to find a framework along the line of OR-cross composition that is known to refute polynomial kernels unless ${\sf NP\subseteq coNP/poly}$.